### v5

# Recent Advances in Stochastic Approximation with Applications to Optimization and Reinforcement Learning

Abstract
Ever since its introduction in the classic paper of Robbins and Monro in 1951, Stochastic Approximation (SA) has become a standard tool for finding a solution of an equation of the fortm ${\bf f}({\boldsymbol{\theta}})={\bf 0}$, when only noisy measurements of ${\bf f}(\cdot)$ are available. In most situations, every component of the putative solution ${\boldsymbol{\theta}}_{t}$ is updated at each step $t$. In some applications such as $Q$-learning, a key technique in Reinforcement Learning (RL), only one component of ${\boldsymbol{\theta}}_{t}$ is updated at each $t$. This is known as asynchronous SA. The topic of study in the present paper is to study Block Asynchronous SA (BASA), in which, at each step $t$, some but not necessarily all components of ${\boldsymbol{\theta}}_{t}$ are updated. The theory presented here embraces both conventional (synchronous) SA as well as asynchronous SA, and all in-between possibilities. We also prove bounds on the rate of convergence of ${\boldsymbol{\theta}}_{t}$ to the solutions.
As a prelude to the new results, we also briefly survey some results on the convergence of the Stochastic Gradient method, proved in a companion papr by the present authors.
2020 Mathematics Subject Classification: 62L20 · 60G17 · 93D05
1 Introduction
1.1 Background
Ever since its introduction in the classic paper of Robbins and Monro [32], Stochastic Approximation (SA) has become a standard tool in many problems in applied mathematics. It is worth noting that the phrase “Stochastic Approximation” was coined in [32], and that the paper does not have a single reference. As stated in [32], the original problem formulation in SA was to find a solution to an equation of the form [^1] For the convenience of the reader, all results cited from the literature are stated in the notation used in the present paper, which may differ from the original paper.
$f(\theta)=c,$
where $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$, $c$ is a specified constant, and one has access only to noisy measurements of the function. Obviously, one can redefine $f$ and assume that $c=0$, without loss of generality. Almost at once, the approach was extended to finding a stationary point of a ${\mathcal{C}}^{1}$-function $J(\cdot):{\mathbb{R}}\rightarrow{\mathbb{R}}$ in [18], and to the case where $J(\cdot):{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ in [3]. Other early contributions are [10, 9]. In the early papers, SA was analyzed under extremely stringent assumptions on the function, and on the measurement error. With the passage of time, subsequent researchers have substantially relaxed the assumptions.
Over the years, SA has become a standard tool for analyzing the behavior of stochastic algorithms in a variety of areas, out of which two topics are the focus in the present paper, namely: optimization, and Reinforcement Learning (RL). The aim of the present paper is two-fold: First, we survey some known results in the theory of SA, including some results due to the present authors. Second, we present some new results on so-called Block, or Batch, SA.
1.2 Problem Formulation
Suppose ${\bf f}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ is some function. It is desired to find a solution to the equation ${\bf f}({\boldsymbol{\theta}}^{*})={\bf 0}$, when only noisy measurements of ${\bf f}(\cdot)$ are available. An iterative approach is adopted to solve this equation. Let $t$ denote the iteration count, and choose the initial guess ${\boldsymbol{\theta}}_{0}$ either in a deterministic or a random fashion. At time (or step) $t+1$, the available measurement is ${\bf f}({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}$, where ${\boldsymbol{\xi}}_{t+1}$ is variously referred to as the measurement error or the “noise.” Both phrases are used interchangeably in this paper. The current guess ${\boldsymbol{\theta}}_{t}$ is updated via the formula
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}+{\boldsymbol{\alpha}}_{t}\circ[{\bf f}({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}],$ (1.1)
where ${\boldsymbol{\alpha}}_{t}\in(0,\infty)^{d}$ is called the step size vector, and $\circ$ denotes the Hadamard product. [^2] Recall that if ${\bf a},{\bf b}$ are vectors of equal dimension, then their Hadamard product ${\bf c}={\bf a}\circ{\bf b}$ is defined by $c_{i}:=a_{i}b_{i}$ for all $i$. If ${\bf g}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ is a map and it is desired to find a fixed point of ${\bf g}(\cdot)$, when we can define ${\bf f}({\boldsymbol{\theta}})={\bf g}({\boldsymbol{\theta}})-{\boldsymbol{\theta}}$. This causes (1.1) to become
${\boldsymbol{\theta}}_{t+1}=({\bf 1}_{d}-{\boldsymbol{\alpha}}_{t})\circ{\boldsymbol{\theta}}_{t}+{\boldsymbol{\alpha}}_{t}\circ[{\bf g}({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}],$ (1.2)
where ${\bf 1}_{d}$ denotes the column vector of $d$ ones. In this case, it is customary to restrict ${\boldsymbol{\alpha}}_{t}$ to belong to $(0,1)^{d}$ instead of $(0,\infty)^{d}$. If $J:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is a $C^{1}$-function, and it is desired to find a stationary point of it, then we can define ${\bf f}({\boldsymbol{\theta}})=-\nabla J({\boldsymbol{\theta}})$, in which case (1.1) becomes
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}+{\boldsymbol{\alpha}}_{t}\circ[-\nabla J({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}].$ (1.3)
The choice ${\bf f}({\boldsymbol{\theta}})=-\nabla J({\boldsymbol{\theta}})$ instead of $\nabla J({\boldsymbol{\theta}})$ is used when the objective is to minimize $J(\cdot)$, and $J(\cdot)$ is convex, at least in a neighborhood of the minimum. If the objective is to maximize $J(\cdot)$, then one would choose ${\bf f}({\boldsymbol{\theta}})=\nabla J({\boldsymbol{\theta}})$.
What is described above is the “core” problem formulation. Several variations are possible, depending on the objective of the analysis, the nature of the of the step size vector, and the nature of the error vector ${\boldsymbol{\xi}}_{t+1}$. Some of the most widely studied variations are described next.
Objectives of the Analysis: Historically, the focus has been on proving that sequence of iterations $\{{\boldsymbol{\theta}}_{t}\}$ converges almost surely to a solution of the equation ${\bf f}({\boldsymbol{\theta}})={\bf 0}$ (or its modification for fixed point and stationarity problems). However, it is possible to weaken the requirement to convergence in the mean. This is the objective in [19] and other subsequent papers. An intermediate possibility is convergence in probability. Suppose ${\boldsymbol{\theta}}_{t}\rightarrow{\boldsymbol{\theta}}^{*}$ in probability, and define
$q(t,\epsilon):=\Pr\{\|{\boldsymbol{\theta}}_{t}-{\boldsymbol{\theta}}^{*}\|_{2}>\epsilon\}.$ (1.4)
Then $q(t,\epsilon)\rightarrow 0$ as $t\rightarrow\infty$ for each $\epsilon>0$. The objective is to derive explicit upper bounds for $q(t,\epsilon)$. Some authors refer to such bounds as “high probability bounds.” The advantage of bounds on $q(t,\epsilon)$ is that they are applicable for all $t$ (or at least, for all sufficiently large $t$), and not just when $t\rightarrow\infty$. For this reason, some authors refer to the derivation of such bounds as finite-time SA. Some contributions in this direction are [41, 35, 2, 7, 30]. We do not discuss FTSA in the paper. The interested reader is referred to the above-cited papers and the references therein.
Step Size Sequences: Next we discuss various options for the step size vector ${\boldsymbol{\alpha}}_{t}$, which is allowed to be random. In all cases, it is assumed that there is a scalar deterministic sequence $\{\beta_{t}\}$ taking values in $(0,\infty)$, or in $(0,1)^{d}$ in the case of (1.2). We will discuss three commonly used variants of SA, namely: synchronous (also called fully synchronous), asynchronous, and block asynchronous. In synchronous SA, one chooses ${\boldsymbol{\alpha}}_{t}=\beta_{t}{\bf 1}_{d}$. Thus, in (1.1), the same step size $\beta_{t}$ is applied to every component of ${\boldsymbol{\theta}}_{t}$. In block asynchronous SA (or BASA), there are $d$ different $\{0,1\}$-valued stochastic processes, denoted by $\kappa_{t}^{i},i\in[d]$, called the “update” processes. Then the $i$-th component of ${\boldsymbol{\theta}}_{t}$ is updated only if $\kappa_{t}^{i}=1$. To put it another way, define the “update set” as
$S_{t}:=\{i\in[d]:\kappa_{t}^{i}=1\}.$
Then $\alpha_{t}^{i}=0$ if $i\not\in S_{t}$. However, this raises the question as to what $\alpha_{t}^{i}$ is for $i\in S_{t}$. Two options are suggested in the literature, known as the “global” clock and the “local” clock respectively. This distinction was first suggested in [4]. If a global clock is used, then $\alpha_{t}^{i}=\beta_{t}$. To define the step size when a local clock is used, first define
$\nu_{t}^{i}:=\sum_{\tau=0}^{t}\kappa_{t}^{i}.$ (1.5)
Thus $\nu_{t}^{i}$ counts the number of times that $\theta_{t}^{i}$ is updated, and is referred to as the “counter” process. Then the step size is defined as
$\alpha_{t}^{i}:=\beta_{\nu_{t}^{i}}.$ (1.6)
The distinction between global and local clocks can be briefly summarized as follows: When a global clock is used, every component of ${\boldsymbol{\theta}}_{t}$ that gets updated has exactly the same step size, namely $\beta_{t}$, while the other components have a step size of zero. When a local clock is used, among the components of ${\boldsymbol{\theta}}_{t}$ that get updated at time $t$, different components may have different step sizes. An important variant of BASA is asynchronous SA (ASA). This phrase was apparently first used in [38], in the context of proving the convergence of the $Q$-learning algorithm from Reinforcement Learning (RL); this algorithm is discussed in Section 5. In ASA, exactly one component of ${\boldsymbol{\theta}}_{t}$ is updated at each $t$. This can be represented as follows: Let $\{N_{t}\}$ be an integer-valued stochastic process taking values in $[d]$. Then, at time $t$, the update set $S_{t}$ is the singleton $\{N_{t}\}$. The counter process $\nu_{t}^{i}$ is now defined via
$\nu_{t}^{i}=\sum_{\tau=0}^{t}I_{\{N_{\tau}=i\}},$
where $I$ denotes the indicator process. The step size can either be $\beta_{t}$ if a global clock is used, or $\beta_{\nu_{t}^{i}}$ if a local clock is used. In [4], the author analyzes the convergence of ASA with both global as well as local clocks. In the $Q$-learning algorithm introduced in [42], the update is asynchronous (one component at a time) and a global clock is used. In [38], where the phrase ASA was first introduced, the convergence of ASA is proved under some assumptions which include $Q$-learning as a special case. Accordingly, the author uses a global clock in the formulation of ASA. In [11], the authors use a local clock to study the rate of convergence of $Q$-learning.
Error Vector: Next we discuss the assumptions made on the error vector ${\boldsymbol{\xi}}_{t+1}$. To state the various assumptions precisely, let ${\boldsymbol{\theta}}_{0}^{t}$ denote $({\boldsymbol{\theta}}_{0},\cdots,{\boldsymbol{\theta}}_{t})$, and define ${\boldsymbol{\alpha}}_{0}^{t}$ and ${\boldsymbol{\xi}}_{1}^{t}$ similarly; note that there is no ${\boldsymbol{\xi}}_{0}$. Let ${\mathcal{F}}_{t}$ denote the $\sigma$-algebra generated by ${\boldsymbol{\theta}}_{0},{\boldsymbol{\alpha}}_{0}^{t},{\boldsymbol{\xi}}_{1}^{t}$, and observe that ${\mathcal{F}}_{t}\subseteq{\mathcal{F}}_{t+1}$. Thus $\{{\mathcal{F}}_{t}\}_{t\geq 0}$ is a filtration; Now (1.1) makes it clear that ${\boldsymbol{\theta}}_{t}$ is measurable with respect to ${\mathcal{F}}_{t}$, denoted by ${\boldsymbol{\theta}}_{t}\in{\mathcal{M}}({\mathcal{F}}_{t})$. Given an ${\mathbb{R}}^{d}$-valued random variable $X$, let $E_{t}(X)$ denote $E(X|{\mathcal{F}}_{t})$, the conditional expectation of $X$ with respect to ${\mathcal{F}}_{t}$, and let $CV_{t}(X)$ denote the conditional variance of $X$, defined as
$CV_{t}(X)=E_{t}(\|X-E_{t}(X)\|_{2}^{2}).$
An important ingredient in SA theory is the set of assumptions imposed on the two entities $E_{t}({\boldsymbol{\xi}}_{t+1})$ and $CV_{t}({\boldsymbol{\xi}}_{t+1})$. We begin with $E_{t}({\boldsymbol{\xi}}_{t+1})$, The simplest assumptions are that
$E_{t}({\boldsymbol{\xi}}_{t+1})={\bf 0},\;\forall t,$ (1.7)
and that there exists a constant $M$ such that
$CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M,\;\forall t.$ (1.8)
where the equality and the bound hold almost surely. To avoid tedious repetition, the phrase “almost surely” is omitted hereafter, unless it is desirable to state it explicitly. Equation (1.7) implies that $\{{\boldsymbol{\xi}}_{t}\}$ is a martingale difference sequence with respect to the filtration $\{{\mathcal{F}}+t\}$. Equation (1.7) firther means that ${\bf f}({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}$ provides an unbiased measurement of ${\bf f}({\boldsymbol{\theta}}_{t})$. In (1.9), the bound on $CV({\boldsymbol{\xi}}_{t+1})$ is not just uniform over $t$, but also uniform over ${\boldsymbol{\theta}}_{t}$. Over time, the assumptions on both $E_{t}({\boldsymbol{\xi}}_{t+1})$ and $CV_{t}({\boldsymbol{\xi}}_{t+1})$ have been relaxed by successive authors. The most general set of conditions to date are found in [17], and are as follows: There exist sequences of constants $\mu_{t}$ and $M_{t}$ such that
$\|E_{t}({\boldsymbol{\xi}}_{t+1})\|_{2}\leq\mu_{t}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}),\;\forall t.$ (1.9)
$CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M_{t}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}^{2}),\;\forall t.$ (1.10)
Convergence conditions for SA under the above assumptions on the error ${\boldsymbol{\xi}}_{t+1}$ are derived in [17], and are stated without proof in Section 2.2 below. The literature review in [17, Section 1.1] contains details of the various intermediate stages between (1.7)–(1.8) and (1.9)–(1.10), and the relevant publications. A condensed version of it is reproduced in Section 2.1. The reader is also directed to [22] for a partial survey that is up to date until its date of publication, 2003.
Methods of Analysis: There are two broad approaches to the analysis of SA, which might be called the ODE approach and the martingale approach. In the ODE approach, it is shown that, as the step sizes $\alpha_{t}\rightarrow 0$, the stochastic sample paths of (1.1) “converge” to the (deterministic) solution trajectories of the associated ODE $\dot{{\boldsymbol{\theta}}}={\bf f}({\boldsymbol{\theta}})$. This approach is introduced in [19, 24, 8]. Book-length treatments of this approach can be found in [20, 21, 1, 6]. The martingale method was pioneered in [13], and independently discovered and enhanced in [31]. In this approach, the stochastic process $\{{\boldsymbol{\theta}}_{t}\}$ s directly analyzed without recourse to any ODE. Conclusions about the behavior of this stochastic process are drawn using the theory of supermartingales. Overall, the proofs are much simpler in the martingale approach. Further, when the function ${\bf f}(\cdot)$ is discontinuous, the martingale approach applies as is. In contrast, when the function ${\bf f}(\cdot)$ is discontinuous, in the ODE approach the limit is not an ODE, but a differential inclusion. Moreover, the typical theorem based on the ODE approach states that if the iterations remain bounded almost surely, then convergence takes place. In contrast, in the martingale approach, both the boundedness and the convergence of the iterations can be established simultaneously. However, the ODE approach is, at least in principle, applicable to the case where ${\bf f}({\boldsymbol{\theta}})={\bf 0}$ has multiple solutions. An important paper in the ODE approach is [5], in which the boundedness of the iterations is a conclusion and not a hypothesis. While this is a substantial contribution, in this approach it is assumed that the equation ${\bf f}({\boldsymbol{\theta}})={\bf 0}$ has a unique solution, thus negating an advantage of the ODE approach. Moreover, it is shown in [40] that if the function ${\bf f}(\cdot)$ grows sublinearly, that is, if
$\lim_{\|{\boldsymbol{\theta}}\|_{2}\rightarrow\infty}\frac{\|{\bf f}({\boldsymbol{\theta}})\|_{2}}{\|{\boldsymbol{\theta}}\|_{2}}=0,$
then the main theorem in [5] does not apply.
1.3 Contributions of the Paper
Leaving aside the survey of the Stochastic Gradient method, the major part of the paper consists of an analysis of Block (or Batch) Asynchronous SA, or BASA, as specified in (1.2). We begin by analyzing the convergence of “intermittently updated” processes of the form
$w_{t+1}=(1-\alpha_{t}\kappa_{t})w_{t}+\alpha_{t}\kappa_{t}\xi_{t+1},$
where $\{w_{t}\}$ is an ${\mathbb{R}}$-valued stochastic process, $\{\xi_{t}\}$ is the measurement error, $\{\alpha_{t}\}$ is a $(0,1)$-valued “step size” process, and $\{\kappa_{t}\}$ is a $\{0,1\}$-valued “update” process. For this formulation, we derive sufficient conditions for convergence, as well as bounds on the rate of convergence. We study both the use of both a local clock as well as a global clock, a distinction first introduced in [4]. This formulation is a precursor to the full BASA formulation of (1.2), where again we derive both sufficient conditions for convergence, and bounds on the rate of convergence. Next, we apply these results to derive sufficient conditions for the convergence of the well-known $Q$-learning algorithm of Reinforcement Learning. Recall that ASA was first proposed in [38], to establish the convergence of the $Q$-learning algorithm using an analog of the Robbins-Monro conditions. However, verifying these conditions is not trivial; indeed, this problem is very similar to an NP-hard problem; see Section 5 for details. In this paper, by introducing a batch version of $Q$-learning, we are able to do away with these unverifiable conditions, and replace them by simple algebraic conditions.
1.4 Scope and Organization of the Paper
This paper contains a survey of some results due to the present authors, and some new results. In Section 2, various results from [17] are stated without proof; these results pertain to the convergence of the synchronous SA algorithm, when the error signal ${\boldsymbol{\xi}}_{t+1}$ satisfies the bounds (1.9) and (1.10). These are the most general assumptions to date. In Section 3, we survey some applications of these convergence results to the stochastic gradient method. Two variants are studied, namely, noise-corrupted exact gradients, and approximate gradients. These two sections comprise the survey part of the paper.
In Section 4, we commence presenting some new results. Specifically, we study Block (or Batch) Asynchronous SA, denoted by BASA, as described in (1.2). The focus is on finding a fixed point of a map ${\bf g}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ which is a contraction in the $\ell_{\infty}$-norm, or a scaled version thereof. This problem arises in Reinforcement Learning in sevdral situations, specifically, $Q$-learning which is studied here in Section 5. The novelties here are that (i) we permit an absolutely general model for choosing the coordinates of ${\boldsymbol{\theta}}_{t}$ to be updated at tim $t$, and (ii) we also derive bounds on the rate of convergence.
2 Synchronous Stochastic Approximation
2.1 Historical Review
We begin with the classical results, starting with [32] which introduced the SA algorithm for the scalar case where $d=1$. However, we state it here for the multidimensional case. In that paper, the update equation is (1.1), and the error ${\boldsymbol{\xi}}_{t+1}$ is assumed to satisfy the following assumptions (though this notation is not used in that paper)
$E_{t}({\boldsymbol{\xi}}_{t+1})={\bf 0},CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M^{2}$ (2.1)
for some finite constant $M$. The first assumption implies that $\{{\boldsymbol{\xi}}_{t+1}\}$ is a martingale difference sequence, and also that ${\bf f}({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1}$ is an unbiased measurement of ${\bf f}({\boldsymbol{\theta}}_{t})$. The second assumption means that the conditional variance of the error is globally bounded, both as a function of ${\boldsymbol{\theta}}_{t}$ and as a function of $t$. With the assumptions in (2.1), along with some assumptions on the function ${\bf f}(\cdot)$, it is shown in [32] that ${\boldsymbol{\theta}}_{t}$ converges to a solution of ${\bf f}({\boldsymbol{\theta}}^{*})={\bf 0}$, provided the step size sequence satisfies the Robbins-Monro (RM) conditions
$\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty,\sum_{t=0}^{\infty}\alpha_{t}=\infty.$ (2.2)
This approach was extended in [18] to finding a stationary point of a ${\mathcal{C}}^{1}$ function $J:{\mathbb{R}}\rightarrow{\mathbb{R}}$, that is, a solution to $\nabla J({\boldsymbol{\theta}})={\bf 0}$,333 Strictly speaking, we should use $J^{\prime}(\theta)$ for the scalar case. But we use vector notation to facilitate comparison with later formulas. using an approximate gradient of $J(\cdot)$. The specific formulation used in [18] is
$h_{t+1}:=\frac{J(\theta_{t}+c_{t}\Delta+\xi_{t+1}^{+})-J(\theta_{t}-c_{t}\Delta+\xi_{t+1}^{-})}{2c_{t}}\approx\nabla J(\theta_{t}).$ (2.3)
where $c_{t}$ is called the increment, $\Delta$ is some fixed number, and $\xi_{t+1}^{+}$, $\xi_{t+1}^{-}$ are the measurement errors. This terminology “increment” is not standard but is used here. In order to make the expression a better and better approximation to the true $\nabla J({\boldsymbol{\theta}}_{t})$, the increment $c_{t}$ must approach zero as $t\rightarrow\infty$. Note that there are two sources of error in (2.3). First, even the errors $\xi^{\pm}_{t+1}$ is zero, the first-order difference is not exactly equal to the gradient $\nabla J({\boldsymbol{\theta}}_{t})$. Second, the presence of the measurement errors $\xi^{\pm}_{t+1}$ introduces an additional error term. To analyze this, let us define
${\bf z}_{t}=E_{t}({\bf h}_{t+1}),{\bf x}_{t}={\bf z}_{t}-\nabla J({\boldsymbol{\theta}}_{t}),{\boldsymbol{\zeta}}_{t+1}={\bf h}_{t+1}-{\bf z}_{t}.$ (2.4)
In this case, the error term satisfies
$\|E_{t}({\boldsymbol{\zeta}}_{t+1})\|_{2}\leq Kc_{t},CV_{t}({\boldsymbol{\zeta}}_{t+1})\leq M^{2}/c_{t}^{2},$ (2.5)
for suitable constants $K,M$. These conditions are more general than in (2.1). In this case, for the scalar case, it was shown in [18] that ${\boldsymbol{\theta}}_{t}$ converges to a stationary point of $J(\cdot)$ if the Kiefer-Wolfwitz-Blum (KWB) conditions
$c_{t}\rightarrow 0,\sum_{t=0}^{\infty}(\alpha_{t}^{2}/c_{t}^{2})<\infty,\sum_{t=0}^{\infty}\alpha_{t}c_{t}<\infty,\sum_{t=0}^{\infty}\alpha_{t}=\infty$ (2.6)
are satisfied. This approach was extended to the multidimensional case in [3], and it is shown that the same conditions also ensure convergence when $d>1$. Note that the conditions automatically imply the finiteness of the sum of $\alpha_{t}^{2}$.
Now we summarize subsequent results. It can be seen from Theorem 2.5 below that in the present paper, the error ${\boldsymbol{\xi}}_{t+1}$ is assumed to satisfy the following assumptions:
$\|E_{t}({\boldsymbol{\xi}}_{t+1})\|_{2}\leq\mu_{t}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}),$ (2.7)
$CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M_{t}^{2}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}^{2}),$ (2.8)
where ${\boldsymbol{\theta}}_{t}$ is the current iteration. It can be seen that the above assumptions extend (2.5) in several ways. First, the conditional expectation is allowed to grow as an affine function of $\|{\boldsymbol{\theta}}_{t}\|_{2}$, for each fixed $t$. Second, the conditional variance is also allowed to grow as a quadratic function of $\|{\boldsymbol{\theta}}_{t}\|_{2}$, for each fixed $t$. Third, while the coefficient $\mu_{t}$ is required to approach zero, the coefficient $M_{t}$ can grow without bound as a function of $t$. We are not aware of any other paper that makes such general assumptions. However, there are several papers wherein the assumptions on ${\boldsymbol{\xi}}_{t+1}$ are intermediate between (2.1) and (2.5). We attempt to summarize a few of them next. For the benefit of the reader, we state the results using the notation of the present paper.
In [19], the author considers a recursion of the form
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}-\alpha_{t}\nabla J({\boldsymbol{\theta}}_{t})+\alpha_{t}{\boldsymbol{\xi}}_{t+1}+\alpha_{t}{\boldsymbol{\beta}}_{t+1},$
where ${\boldsymbol{\beta}}_{t}\rightarrow{\bf 0}$ as $t\rightarrow\infty$. Here, the sequence $\{{\boldsymbol{\xi}}_{t+1}\}$ is not assumed to be a martingale difference sequence. Rather, it is assumed to satisfy a different set of conditions, referred to as the Kushner-Clark conditions; see [19, A5]. It is then shown that if the error sequence $\{{\boldsymbol{\xi}}_{t+1}\}$ satisfies (2.1), i.e., is a martingale difference sequence, then Assumption (A5) holds. Essentially the same formulation is studied in [25]. The same formulation is also studied [6, Section 2.2], where (2.1) holds, and ${\boldsymbol{\beta}}_{t}\rightarrow{\bf 0}$ as $t\rightarrow\infty$. In [37], it is assumed only that $\limsup_{t}{\boldsymbol{\beta}}_{t}<\infty$. In all cases, it is shown that ${\boldsymbol{\theta}}_{t}$ converges to a solution of ${\bf f}({\boldsymbol{\theta}}^{*})={\bf 0}$, provided the iterations remain bounded almost surely. However, the assumption that $\beta_{t}\rightarrow 0$ as $t\rightarrow\infty$ may not, by itself, be sufficient to ensure that the iterations are bounded. Ideally, the boundedness of the iterations should be a conclusion and not a hypothesis.
In all of the above references, the bound on $CV_{t}({\boldsymbol{\xi}}_{t+1})$ is as in (2.1). We are aware of only one paper when the bound on $CV_{t}({\boldsymbol{\xi}}_{t+1})$ is akin to that in (2.8). In [15], the authors study smooth convex optimization. They assume that the estimated gradient is unbiased, so that $\mu_{t}=0$ for all $t$. However, an analog of (2.8) is assumed to hold, which is referred to as “state-dependent noise.” See [15, Assumption (SN)]. In short, there is no paper wherein the assumptions on the error are as general as in (2.7) and (2.8).
2.2 Convergence Theorems
In this subsection, we state without proof some results from [17] on the convergence of SA, when the measurement error satisfies (2.7) and (2.8), which are the most general assumptions to date. In addition to proving convergence, we also provide a general framework for estimating the rate of convergence. The applications of these convergence theorems to stochastic gradient descent (SGD) are discussed in Section 3.
The theorems proved in [17] make use of the following classic “almost supermartingale theorem” of Robbins-Siegmund [31, Theorem 1]. The result is also proved as [1, Lemma 2, Section 5.2]. Also see a recent survey paper as [12, Lemma 4.1]. The theorem states the following:
Lemma 2.1.
Suppose $\{z_{t}\},\{f_{t}\},\{g_{t}\},\{h_{t}\}$ are stochastic processes taking values in $[0,\infty)$, adapted to some filtration $\{{\mathcal{F}}_{t}\}$, satisfying
$E_{t}(z_{t+1})\leq(1+f_{t})z_{t}+g_{t}-h_{t}\mbox{ a.s.},\;\forall t,$ (2.9)
where, as before, $E_{t}(z_{t+1})$ is a shorthand for $E(z_{t+1}|{\mathcal{F}}_{t})$. Then, on the set
$\Omega_{0}:=\{\omega:\sum_{t=0}^{\infty}f_{t}(\omega)<\infty\}\cap\{\omega:\sum_{t=0}^{\infty}g_{t}(\omega)<\infty\},$
we have that $\lim_{t\rightarrow\infty}z_{t}$ exists, and in addition, $\sum_{t=0}^{\infty}h_{t}(\omega)<\infty$. In particular, if $P(\Omega_{0})=1$, then $\{z_{t}\}$ is bounded almost surely, and $\sum_{t=0}^{\infty}h_{t}(\omega)<\infty$ almost surely.
The first convergence result, namely Theorem 2.4 below, is a fairly straight-forward, but useful, extension of Lemma 2.1. It is based on a concept which is introduced in [13] but without giving it a name. The formal definition is given in [40, Definition 1]:
Definition 2.2.
A function $\eta:{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}_{+}$ is said to belong to Class ${\mathcal{B}}$ if $\eta(0)=0$, and in addition
$\inf_{\epsilon\leq r\leq M}\eta(r)>0,\;\forall 0<\epsilon<M<\infty.$
Note $\eta(\cdot)$ is not assumed to be monotonic, or even to be continuous. However, if $\eta:{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}_{+}$ is continuous, then $\eta(\cdot)$ belongs to Class ${\mathcal{B}}$ if and only if (i) $\eta(0)=0$, and (ii) $\eta(r)>0$ for all $r>0$. Such a function is called a “class P function” in [14]. Thus a Class ${\mathcal{B}}$ function is slightly more general than a function of Class $P$.
As example of a function of Class ${\mathcal{B}}$ is given next:
Example 2.3.
Define a function $f:{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}$ by
$\phi(\theta)=\left\{\begin{array}[]{ll}\theta,&\mbox{if }\theta\in[0,1],\\ e^{-(\theta-1)},&\mbox{if }\theta>1.\end{array}\right.$
Then $\phi$ belongs to Class ${\mathcal{B}}$. A sketch of the function $\phi(\cdot)$ is given in Figure 1. Note that, if we were to change the definition to:
$\phi(\theta)=\left\{\begin{array}[]{ll}\theta,&\mbox{if }\theta\in[0,1],\\ 2e^{-(\theta-1)},&\mbox{if }\theta>1,\end{array}\right.$
then $\phi(\cdot)$ would be discontinuous at $\theta=1$, but it would still belong to Class ${\mathcal{B}}$. Thus a function need not be continuous to belong to Class ${\mathcal{B}}$.
Figure 1: An illustration of a function in Class ${\mathcal{B}}$
Now we present our first convergence theorem, which is an extension of Lemma 2.1. This theorem is used to establish the convergence of stochastic gradient methods for nonconvex functions, as discussed in Section 3. It is [17, Theorem 1].
Theorem 2.4.
Suppose $\{z_{t}\},\{f_{t}\},\{g_{t}\},\{h_{t}\},\{\alpha_{t}\}$ are $[0,\infty)$-valued stochastic processes defined on some probability space $(\Omega,\Sigma,P)$, and adapted to some filtration $\{{\mathcal{F}}_{t}\}$. Suppose further that
$E_{t}(z_{t+1})\leq(1+f_{t})z_{t}+g_{t}-\alpha_{t}h_{t}\mbox{ a.s.},\;\forall t.$ (2.10)
Define
$\Omega_{0}:=\{\omega\in\Omega:\sum_{t=0}^{\infty}f_{t}(\omega)<\infty\mbox{ and }\sum_{t=0}^{\infty}g_{t}(\omega)<\infty\mbox{ and }\sum_{t=0}^{\infty}\alpha_{t}(\omega)=\infty\}.$ (2.11)
Then
  1. 1.
    There exists a random variable $W$ defined on $(\Omega,\Sigma,P)$ such that $z_{t}(\omega)\rightarrow W(\omega)$ for all $\omega\in\Omega_{0}$. Moreover
    $\liminf_{t\rightarrow\infty}h_{t}(\omega)=0\;\forall\omega\in\Omega_{0}.$ (2.12)
  2. 2.
    Suppose there exists a function $\eta(\cdot)$ of Class ${\mathcal{B}}$ such that $h_{t}(\omega)\geq\eta(z_{t}(\omega))$ for all $\omega\in\Omega_{0}$. Then $z_{t}(\omega)\rightarrow 0$ as $t\rightarrow\infty$ for all $\omega\in\Omega_{0}$.
Next we study a linear stochastic recurrence relation. Despite its simplicity, it is a key tool in establishing the convergence of Stochastic Gradient Descent (SGD) studied in Section 3. Suppose ${\boldsymbol{\theta}}_{0}$ is an ${\mathbb{R}}^{d}$-valued random variable, and that $\{{\boldsymbol{\zeta}}_{t}\}_{t\geq 1}$ is an ${\mathbb{R}}^{d}$-valued stochastic process. Define $\{{\boldsymbol{\theta}}_{t},t\geq 1\}$ recursively by
${\boldsymbol{\theta}}_{t+1}=(1-\alpha_{t}){\boldsymbol{\theta}}_{t}+\alpha_{t}{\boldsymbol{\zeta}}_{t+1},t\geq 0,$ (2.13)
where $\{\alpha_{t}\}_{t\geq 0}$ is another $[0,1)$-valued stochastic process. Define $\{{\mathcal{F}}_{t}\}$ to be the filtration where ${\mathcal{F}}_{t}$ is the $\sigma$-algebra generated by ${\boldsymbol{\theta}}_{0},\alpha_{0}^{t},{\boldsymbol{\zeta}}_{1}^{t}$. Note that (2.13) is of the form (1.2) with ${\bf g}({\boldsymbol{\theta}})\equiv{\bf 0}$. Hence ${\bf g}(\cdot)$ has the unique fixed point ${\bf 0}$, and we would want that ${\boldsymbol{\theta}}_{t}\rightarrow{\bf 0}$ as $t\rightarrow\infty$. The next theorem is [17, Theorem 3].
Theorem 2.5.
Suppose there exist sequences of constants $\{\mu_{t}\}$, $\{M_{t}\}$ such that, for all $t\geq 0$ we have
$\|E_{t}({\boldsymbol{\zeta}}_{t+1})\|_{2}=\|{\boldsymbol{\eta}}_{t}\|_{2}\leq\mu_{t}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}),$ (2.14)
$CV_{t}({\boldsymbol{\zeta}}_{t+1})=E_{t}(\|{\boldsymbol{\psi}}_{t+1}\|_{2}^{2})\leq M_{t}^{2}(1+\|{\boldsymbol{\theta}}_{t}\|_{2}^{2}).$ (2.15)
Under these conditions, if
$\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty,\sum_{t=0}^{\infty}\mu_{t}\alpha_{t}<\infty,\sum_{t=0}^{\infty}M^{2}_{t}\alpha^{2}_{t}<\infty,$ (2.16)
then $\{{\boldsymbol{\theta}}_{t}\}$ is bounded, and $\|{\boldsymbol{\theta}}_{t}\|_{2}$ converges to an ${\mathbb{R}}^{d}$-valued random variable. If in addition,
$\sum_{t=0}^{\infty}\alpha_{t}=\infty,$ (2.17)
then ${\boldsymbol{\theta}}_{t}\rightarrow{\bf 0}$.
Next, we state an extension of Theorem 2.4 that provides an estimate on rates of convergence. For the purposes of this paper, we use the following definition inspired by [23].
Definition 2.6.
Suppose $\{Y_{t}\}$ is a stochastic process, and $\{f_{t}\}$ is a sequence of positive numbers. We say that
  1. 1.
    $Y_{t}=O(f_{t})$ if $\{Y_{t}/f_{t}\}$ is bounded almost surely.
  2. 2.
    $Y_{t}=\Omega(f_{t})$ if $Y_{t}$ is positive almost surely, and $\{f_{t}/Y_{t}\}$ is bounded almost surely.
  3. 3.
    $Y_{t}=\Theta(f_{t})$ if $Y_{t}$ is both $O(f_{t})$ and $\Omega(f_{t})$.
  4. 4.
    $Y_{t}=o(f_{t})$ if $Y_{t}/f_{t}\rightarrow 0$ almost surely as $t\rightarrow\infty$.
The next theorem is a modification of Theorem 2.4 that provides bounds on the rate of convergence. It is [17, Theorem 2].
Theorem 2.7.
Suppose $\{z_{t}\},\{f_{t}\},\{g_{t}\},\{\alpha_{t}\}$ are stochastic processes defined on some probability space $(\Omega,\Sigma,P)$, taking values in $[0,\infty)$, adapted to some filtration $\{{\mathcal{F}}_{t}\}$. Suppose further that
$E_{t}(z_{t+1})\leq(1+f_{t})z_{t}+g_{t}-\alpha_{t}z_{t}\;\forall t,$ (2.18)
where
$\sum_{t=0}^{\infty}f_{t}(\omega)<\infty,\sum_{t=0}^{\infty}g_{t}(\omega)<\infty,\sum_{t=0}^{\infty}\alpha_{t}(\omega)=\infty.$
Then $z_{t}=o(t^{-\lambda})$ for every $\lambda\in(0,1]$ such that (i) there exists a $T<\infty$ such that
$\alpha_{t}(\omega)-\lambda t^{-1}\geq 0\;\forall t\geq T,$ (2.19)
and in addition (ii)
$\sum_{t=0}^{\infty}(t+1)^{\lambda}g_{t}(\omega)<\infty,\sum_{t=0}^{\infty}[\alpha_{t}(\omega)-\lambda t^{-1}]=\infty.$ (2.20)
With this motivation, we present a refinement of Theorem 2.5. It is [17, Theorem 4].
Theorem 2.8.
Let various symbols be as in Theorem 2.5. Further, suppose there exist constants $\gamma>0$ and $\delta\geq 0$ such that [^4] Since $t^{-\gamma}$ is undefined when $t=0$, we really mean $(t+1)^{-\gamma}$. The same applies elsewhere also.
$\mu_{t}=O(t^{-\gamma}),M_{t}=O(t^{\delta}),$
where we take $\gamma=1$ if $\mu_{t}=0$ for all sufficiently large $t$, and $\delta=0$ if $M_{t}$ is bounded. Choose the step-size sequence $\{\alpha_{t}\}$ as $O(t^{-(1-\phi)})$ and $\Omega(t^{-(1-c)})$ where $\phi$ is chosen to satisfy
$0<\phi<\min\{0.5-\delta,\gamma\},$
and $c\in(0,\phi]$. Define
$\nu:=\min\{1-2(\phi+\delta),\gamma-\phi\}.$ (2.21)
Then $\|{\boldsymbol{\theta}}_{t}\|_{2}^{2}=o(t^{-\lambda})$ for every $\lambda\in(0,\nu)$. In particular, if $\mu_{t}=0$ for all $t$ and $M_{t}$ is bounded with respect to $t$, then we can take $\nu=1-2\phi$.
3 Applications to Optimization
In this section, we reprise some relevant results from [17] on the convergence of the Stochastic Gradient Method. Two variants are studied, namely: true gradient corrupted by an additive noise (Section 3.2) and approximate gradient which is also corrupted by additive noise (Section 3.3).
3.1 Assumptions
Next, we state various assumptions on the function $J(\cdot)$ that we seek to minimize. Note that it is not assumed that $J(\cdot)$ is convex. First there are a couple of standard assumptions:
  1. (S1).
    The function $J(\cdot)$ is bounded below and has a unique global minimum. By translating coordinates, it is assumed that the global minimum value of $J(\cdot)$ is zero, and occurs at ${\boldsymbol{\theta}}^{*}={\bf 0}$.
  2. (S2.)
    The function $J(\cdot)$ has compact level sets; thus, for each constant $c$, the set
    $L_{J}(c):=\{{\boldsymbol{\theta}}:J({\boldsymbol{\theta}})\leq c\}$
    is compact. Note that some authors use the adjective “coercive” to refer to such functions.
These assumptions are made in every theorem, and thus not repeated every time.
The above assumptions are fairly standard. Now we state the “operative” assumptions. Note that different subsets of assumptions are used in different theorems.
  1. (J1.)
    $J:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$ is ${\mathcal{C}}^{1}$, and $\nabla J(\cdot)$ is Lipschitz-continuous with constant $L$.
  2. (J2.)
    There exists a constant $H$ such that
    $\|\nabla J({\boldsymbol{\theta}})\|_{2}^{2}\leq HJ({\boldsymbol{\theta}}),\;\forall{\boldsymbol{\theta}}\in{\mathbb{R}}^{d}.$ (3.1)
  3. (J3.)
    There exists a function $\psi(\cdot)$ of Class ${\mathcal{B}}$ such that
    $\|\nabla J({\boldsymbol{\theta}})\|_{2}^{2}\geq\psi(J({\boldsymbol{\theta}})),\;\forall{\boldsymbol{\theta}}\in{\mathbb{R}}^{d}.$ (3.2)
  4. (J3’.)
    There exists a constant $K$ such that
    $\|\nabla J({\boldsymbol{\theta}})\|_{2}^{2}\geq KJ({\boldsymbol{\theta}}),\;\forall{\boldsymbol{\theta}}\in{\mathbb{R}}^{d}.$ (3.3)
Now we discuss the significance of these assumptions. Assumption (J1) guarantees that $J({\boldsymbol{\theta}})=O(\|{\boldsymbol{\theta}}\|_{2}^{2})$. If $J(\cdot)$ is convex, then it is known that Assumption (J2) is satisfied with $H=2L$; see [26, Theorem 2.1.5]. If in addition $J(\cdot)$ is also $R$-strongly convex, in the sense that
$J({\boldsymbol{\psi}})\geq J({\boldsymbol{\theta}})+\langle\nabla J({\boldsymbol{\theta}}),{\boldsymbol{\psi}}-{\boldsymbol{\theta}}\rangle+\frac{R}{2}\|{\boldsymbol{\psi}}-{\boldsymbol{\theta}}\|_{2}^{2},$
for some constant $R$, then it follows from [26, Theorem 2.1.10] that
$J({\boldsymbol{\theta}})\leq\frac{1}{2R}\|\nabla J({\boldsymbol{\theta}}\|_{2}^{2},\;\forall{\boldsymbol{\theta}}\in{\mathbb{R}}^{d}.$
Thus $J(\cdot)$ satisfies Assumption (J3’) with $K=1/2R$. Note that Assumption (J3) is weaker than Assumption (J3’). To ilustrate this, consider the map $J:{\mathbb{R}}\rightarrow{\mathbb{R}}$ defined by $J(\theta)=|\theta|^{p}$, where $p\in(1,2)$. Also, there exist nonconvex functions that also satisfy these assumptions. See Examples 3.1 and 3.2 below.
Example 3.1.
Let us return to the function ϕ():++:italic-$\phi(\cdot):{\mathbb{R}}_{+}\rightarrow{\mathbb{R}}_{+}$ defined in Example 2.3. Define $f:{\mathbb{R}}\rightarrow{\mathbb{R}}$ as the odd extension of $\phi$ and $J:{\mathbb{R}}\rightarrow{\mathbb{R}}$ as
[…]
so that $f(\cdot)$ is the derivative of $J(\cdot)$. Figure 2 shows the function $J(\cdot)$ and its derivative $f(\cdot)$.
Refer to caption Refer to caption
The function $f(\cdot)$ The function $J(\cdot)$
Figure 2: A nonconvex function
It is obvious that $J(\cdot)$ is not convex. Moreover, $f(\theta)\rightarrow 0$ as $|\theta|\rightarrow\infty$. Hence $J(\cdot)$ satisfies Assumptions (J1), (J2) and (J3), but not (J3’). Hence SGD converges in this case as a consequence of Theorem 3.3, but Theorem 3.4 does not apply.
Example 3.2.
In this example, we modify the function in Example 2.3 as follows: Define $\psi(0)=0$. If $n$ is an even integer (including zero), then
$\psi(n+t)=\psi(n)+t,t\in[0,1].$
If $n$ is an odd integer, then
$\psi(n+t)=\psi(n)-\exp(t),t\in[0,1].$
Then $\psi(\theta)$ is bounded above and below by linear functions of $\theta$. Hence, if we define $f$ to be the odd extension of $\psi$, and $J$ to be the function with $J(0)=0$ with derivative $f$, then $J$ is not convex. However, it does satisfy Assumptions (J1) , (J2) and (J3’). Hence, by choosing a step-size sequence $\alpha_{t}=O(t^{-(1-\epsilon)})$ $\alpha_{t}=O(t^{-(1-C)})$ where $0<C\leq\epsilon$, we can achieve a convergence rate for the SGD with rate arbitrarily close to $O(t^{-(1-\epsilon)})$.
3.2 Convergence of SGD with Noisy Gradients
The general class of algorithms studied in this subsection and the next is now described. Begin with an initial guess ${\boldsymbol{\theta}}_{0}\in{\mathbb{R}}^{d}$ which could be random. At each time $t+1\geq 1$, choose a search direction ${\bf h}_{t+1}$ which is random; different algorithms differ in the manner in which the search direction is chosen. Choose a step size $\alpha_{t}$, which could also be random. Then update the guess via
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}-\alpha_{t}{\bf h}_{t+1}.$ (3.4)
In the case of the SGD (Stochastic Gradient Descent) algorithm studied in this subsection, we choose
${\bf h}_{t+1}=\nabla J({\boldsymbol{\theta}}_{t})+{\boldsymbol{\xi}}_{t+1},$ (3.5)
where ${\boldsymbol{\xi}}_{t+1}$ is a measurement error. In the next subsection, more general choices for ${\bf h}_{t+1}$ are permitted. Define
${\bf z}_{t}:=E_{t}({\bf h}_{t+1}),{\boldsymbol{\zeta}}_{t+1}:={\bf h}_{t+1}-{\bf z}_{t},$ (3.6)
If the direction ${\bf h}_{t+1}$ is chosen as in (3.5), then
$CV_{t}({\bf h}_{t+1})=E_{t}(\|{\boldsymbol{\xi}}_{t+1}\|_{2}^{2}.$
${\bf z}_{t}=\nabla J({\boldsymbol{\theta}}_{t}),{\boldsymbol{\zeta}}_{t+1}={\boldsymbol{\xi}}_{t+1}.$
Thus (3.5) represents full-coordinate gradient descent using noise-corrupted gradients. The following are the assumptions on the measurement error.
  1. (N1.)
    We have that
    $E_{t}({\boldsymbol{\xi}}_{t+1})={\bf 0},\;\forall t\geq 0.$ (3.7)
    This implies that ${\bf h}_{t+1}$ in (3.5) is an unbiased measurement of $\nabla J({\boldsymbol{\theta}}_{t})$.
  2. (N2.)
    There exists a constant $M$ such that
    $CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M^{2}(1+\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}),\;\forall t\geq 0.$ (3.8)
    Note that if (J2) holds, then
    $CV_{t}({\boldsymbol{\xi}}_{t+1})\leq M^{2}(1+HJ({\boldsymbol{\theta}}_{t})).$ (3.9)
With this background, we state the first convergence result, which does not have any conclusions about the rate of convergence. Subsequently, we replace Assumption (J3) by (J3’), and also derive bounds on the rate of convergence. The next result is [17, Theorem 5].
Theorem 3.3.
Suppose the objective function $J(\cdot)$ satisfies Assumptions (J1) and (J2), and that the error term ${\boldsymbol{\xi}}_{t+1}$ satisfies Assumptions (N1) and (N2). Under these conditions, we have the following conclusions:
  1. 1.
    If
    $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty,$ (3.10)
    then $\{\nabla J({\boldsymbol{\theta}}_{t})\}$ and $\{J({\boldsymbol{\theta}}_{t})\}$ are bounded.
  2. 2.
    If in addition $J(\cdot)$ also satisfies Assumption (J3), and
    $\sum_{t=0}^{\infty}\alpha_{t}=\infty,$ (3.11)
    then $J({\boldsymbol{\theta}})\rightarrow 0$ and $\nabla J({\boldsymbol{\theta}}_{t})\rightarrow{\bf 0}$ as $t\rightarrow\infty$.
Next we strengthen Assumption (J3) to (J3’), and prove an estimate for the rate of convergence. This is [17, Theorem 6].
Theorem 3.4.
Suppose that, in Theorem 3.3, Item 2, Assumption (J3) is replaced by Assumption (J3’), and that the step size sequence satisfies
$\alpha_{t}=O(t^{-(1-\phi)}),\alpha_{t}=\Omega(t^{-(1-c)}),c\in(0,\phi]$
for some $\phi\in(0,0.5)$. Then $J({\boldsymbol{\theta}}_{t}),\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}=o(t^{-\lambda})$ for every $\lambda<1-2\phi$.
Next we discuss an “optimal” choice for $\phi$. Clearly, the smaller $\phi$ is, the closer $\lambda$ is to $1$. Hence, if we choose $\phi=\epsilon$, a very small number, then $J({\boldsymbol{\theta}}_{t}),\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}=o(t^{-\lambda})$ for $\lambda$ arbitrarily close to $1-2\epsilon$. The conclusion is that the bound is optimized with very very small values of $\phi$, or step sizes decreasing just slightly faster than $(t+1)^{-1}$. With such a choice, $J({\boldsymbol{\theta}}_{t})=o(t^{-\lambda})$ for $\lambda$ arbitrarily close to $1$. Recall that $O(t^{-1})$ is the rate of convergence for the steepest descent method with a strongly convex objective function; see [27, Corollary 2.1.2] and [27, Eq. (2.1.39)]. Thus an important consequence of Theorem 3.4 is that, even with noisy measurements of the gradient, SGD can be made to converge at a rate arbitrarily close to the best achievable rate with noise-free measurements.
3.3 Convergence of SGD with Approximate Gradients
In this section, we analyze the convergence of the SGD method when the exact gradient is replaced by a first-order approximation. Specifically, we use the so-called SPSA (Simultaneous Perturbation Stochastic Approximation) approach first presented in [34, 33]; see also [28] for related ideas.
Recall the general form of the optimization algorithm:
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}-\alpha_{t}{\bf h}_{t+1}.$ (3.12)
Now define
${\bf z}_{t}=E_{t}({\bf h}_{t+1}),{\boldsymbol{\zeta}}_{t+1}={\bf h}_{t+1}-{\bf z}_{t},{\bf x}_{t}={\bf z}_{t}-\nabla J({\boldsymbol{\theta}}_{t}).$ (3.13)
In the previous subsection, we had ${\bf z}_{t}=\nabla J({\boldsymbol{\theta}}_{t})$, so that ${\bf h}_{t+1}$ is a noisy version of the gradient at ${\boldsymbol{\theta}}_{t}$. In this subsection, we replace the true gradient by a first-order approximation, as defined next. This approach, introduced in [33], requires only two function evaluations at each $t$, irrespective of the value of $d$. Define
$h_{t+1,i}^{(B)}=\frac{[J({\boldsymbol{\theta}}_{t}+c_{t}{\boldsymbol{\Delta}}_{t+1})+\xi_{t+1,i}^{+}]-[J({\boldsymbol{\theta}}_{t}-c_{t}{\boldsymbol{\Delta}}_{t+1})-\xi_{t+1,i}^{-}]}{2c_{t}\Delta_{t+1,i}},$ (3.14)
where $\Delta_{t+1,i},i\in[d]$ are $d$ different and pairwise independent Rademacher variables. [^5] Recall that Rademacher random variables assume values in $\{-1,1\}$ and are independent of each other. Moreover, it is assumed that $\Delta_{t+1,i},i\in[d]$ are all independent of the filtration ${\mathcal{F}}_{t}$. The measurement errors $\xi_{t+1,1}^{+},\cdots,\xi_{t+1,d}^{+}$, $\xi_{t+1,1}^{-},\cdots,\xi_{t+1,d}^{-}$ represent the measurement errors. It is assumed that these $2d$ random variables are pairwise independent with zero conditional expectation with respect to ${\mathcal{F}}_{t}$, and are also independent of $\Delta_{t+1,j}$ for each $j\in[d]$. In [28], instead of Rademacher variables, a random Gaussian vector is used. The advantage of using Rademacher variables is that the ratio $\|{\boldsymbol{\Delta}}_{t+1}\|_{2}/|\Delta_{t+1,i}|$ always equals $\sqrt{d}$.
If we define
$y_{t+1,i}^{(B)}:=\frac{J({\boldsymbol{\theta}}_{t}+c_{t}{\boldsymbol{\Delta}}_{t+1})-J({\boldsymbol{\theta}}_{t}-c_{t}{\boldsymbol{\Delta}}_{t+1})}{2c_{t}\Delta_{t+1,i}},$ (3.15)
$\xi_{t+1,i}^{(B)}:=\frac{\xi_{t+1,i}^{+}-\xi_{t+1,i}^{-}}{2c_{t}\Delta_{t+1,i}},$ (3.16)
then one can express
${\bf h}_{t+1}^{(B)}={\bf y}_{t+1}^{(B)}+{\boldsymbol{\xi}}_{t+1}^{(B)}.$ (3.17)
However, in general it not true that ${\bf y}_{t+1}^{(B)}$ is measurable with respect to ${\mathcal{F}}_{t}$, though it continues to be true that $E_{t}({\boldsymbol{\xi}}_{t+1}^{(B)})={\bf 0}$. Therefore
${\bf z}_{t}^{(B)}=E_{t}({\bf y}_{t+1}^{(B)}),{\boldsymbol{\zeta}}_{t+1}^{(B)}={\bf y}_{t+1}^{(B)}-{\bf z}_{t}^{(B)}+{\boldsymbol{\xi}}_{t+1}^{(B)}.$ (3.18)
The next result is [17, Theorem 8].
Theorem 3.5.
Suppose there exists a constant $M$ such that
$CV_{t}({\boldsymbol{\zeta}}_{t+1})\leq M(1+\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}).$
  1. 1.
    Suppose Assumptions (J1), (J2) and (J3) hold. Then the iterations of the Stochastic Gradient Descent algorithm are bounded almost surely whenever
    $\sum_{t=0}^{\infty}\alpha_{t}^{2}<\infty,\sum_{t=0}^{\infty}\alpha_{t}c_{t}<\infty,\sum_{t=0}^{\infty}\alpha_{t}^{2}/c_{t}^{2}<\infty,$ (3.19)
  2. 2.
    If, in addition to (3.19), we also have
    $\sum_{t=0}^{\infty}\alpha_{t}=\infty,$ (3.20)
    then $J({\boldsymbol{\theta}}_{t})\rightarrow 0$ and $\nabla J({\boldsymbol{\theta}}_{t})\rightarrow{\bf 0}$ almost surely as $t\rightarrow\infty$.
  3. 3.
    Suppose Assumption (J3) is strengthened to Assumption (J3’). Further, suppose that $\alpha_{t}=O(t^{-(1-\phi)})$ and $\alpha_{t}=\Omega(t^{-(1-C)})$, where $C\in(0,\phi]$, and $c_{t}=\Theta(t^{-s})$. Suppose further that
    $\phi<s,\phi+s<0.5,$ (3.21)
    and define
    $\nu:=\min\{1-2(\phi+s),s-\phi\}.$ (3.22)
    Then
    $J({\boldsymbol{\theta}}_{t}),\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}=O(t^{-\lambda})\;\forall\lambda<\nu.$ (3.23)
The reader will recognize that (3.19) and (3.20) are the Kiefer-Wolfowitz-Blum conditions mentioned earlier.
We conclude this section by determing “optimal” choices for the step size exponent $\phi$ and the increment exponent $s$ so as to the get the best possible bound for $\lambda$. The black lines in Figure 1 show the region in the $\phi$-$s$ plane where (3.21) is satisfied. Now let us study (3.22), and set the two parts in the minimum operator equal, that is
$11-2(\phi+s)=s-\phi,\mbox{ or }3s+\phi=$ 1 .
This is the blue line where $\nu$ equals either of the two terms in (3.22). It is obvious that the largest value for $\nu$ over the feasible region, and hence the largest bound for the rate of convergence $\lambda$, occurs when $\phi=0,s=1/3$. Strictly speaking this is not a part of the feasible region. So the conclusion is to choose the increment sequence to be $c_{t}=\Theta(t^{-(1/3-\epsilon)})$, and the step size sequence $\alpha_{t}=\Theta(t^{-(1-\epsilon)})$ for a small value of $\epsilon$. This leads to the “optimal” rate of convergence for $J({\boldsymbol{\theta}}_{t})$ and $\|\nabla J({\boldsymbol{\theta}}_{t})\|_{2}^{2}$, namely $o(t^{-(1/3-\epsilon)})$. As mentioned above, this can be compared with the rate of $O(t^{-1/2})$ when approximate gradients are used with noise-free function measurements; see [28]. Of course, it is possible that future research might lead to better bounds.
$\phi$$s$1/2121/21 / 21/3131/31 / 31/4141/41 / $41/$ 4
Figure 3: Feasible combinations of $(\phi,s)$
4 Block Asynchronous Stochastic Approximation
Until now, we have reviewed some results from a companion paper [17]. This section and the next contain original results due to the authors that are not reported anywhere else. Suppose one wishes to solve (1.2), that is, to find a fixed point of a given map ${\bf g}(\cdot)$. As mentioned earlier, when every component of ${\boldsymbol{\theta}}_{t}$ is updated at each $t$, this is the standard version of SA, referred to by us “synchronous” SA, though the term is not very standard. When exactly one component of ${\boldsymbol{\theta}}_{t}$ is updated at each $t$, this is known as “Asynchronous” SA, a term first introduced in [38]. In this section, we study the solution of (1.2) using “Block Asynchronous” SA, whereby, At each step $t$, some but not necessarily all components of ${\boldsymbol{\theta}}_{t}$ are updated. This is denoted by the acronym BASA. Clearly both Synchronous SA and Asynchronous SA are special cases of BASA.
4.1 Intermittent Updating: Convergence and Rates
The key distinguishing feature of BASA is that each component of ${\boldsymbol{\theta}}_{t}$ gets updated in an “intermittent” fashion. Before tackling the convergence of BASA in full generality, in the present subsection we state and prove results analogous to Theorems 2.5 and 2.8 for the scalar case with intermittent updating.
The problem setup is as follows: The recurrence relationship is
$w_{t+1}=(1-\alpha_{t}\kappa_{t})w_{t}+\alpha_{t}\kappa_{t}\xi_{t+1},$ (4.1)
where $\{w_{t}\}$ is an ${\mathbb{R}}$-valued stochastic process of interest, $\{\xi_{t}\}$ is the measurement error (or “noise”), $\{\alpha_{t}\}$ is a $(0,1)$-valued stochastic process called the “step size” process, and $\{\kappa_{t}\}$ is a $\{0,1\}$-valued stochastic process called the “update” process. Clearly, if $\kappa_{t}=0$, then $w_{t+1}=w_{t}$, irrespective of the value of $\alpha_{t}$; therefore $w_{t+1}$ is updated only at those $t$ for which $\kappa_{t}=1$. This is the rationale for the name. With the update process $\{\kappa_{t}\}$, as before we associate a “counter” process $\{\nu_{t}\}$, defined by
$\nu_{t}=\sum_{s=0}^{t}\kappa_{s}.$ (4.2)
Thus $\nu_{t}$ is the number of times up to and including time $t$ at which $w_{t}$ is updated. We also define
$\nu^{-1}(\tau):=\min\{t\in{\mathbb{N}}:\nu_{t}=\tau\},\;\forall\tau\geq 1.$ (4.3)
Then $\nu^{-1}(\cdot)$ is well-defined, and
$\nu(\nu^{-1}(\tau))=\tau,\nu^{-1}(\nu_{t})\leq t,\nu^{-1}(\tau)\leq\tau-1.$ (4.4)
The last inequality arises from the fact that there are $t+1$ terms in (4.2). Also, $\kappa_{t}=1$ only when $t=\nu^{-1}(\tau)$ for some $\tau$, and is zero for other values of $t$. Hence, in (4.1), if $t=\nu^{-1}(\tau)$ for some $\tau$, then $w_{t}$ gets updated to $w_{t+1}$, and
$w_{t+1}=w_{t+2}=\cdots=w_{\nu^{-1}(\tau+1)},$ (4.5)
at which time $w$ gets updated again. Thus $w_{t}$ is a “piecewise-constant” process, remaining constant between updates. This suggests that we can transform the independent variable from $t$ to $\tau$. Define
$x_{\tau}:=w_{\nu^{-1}(\tau)},\zeta_{\tau+1}:=\xi_{\nu^{-1}(\tau)+1},\;\forall\tau\geq 1,$ (4.6)
with the convention that $x_{1}=w_{0}$. Note that the convention is consistent whether $\nu_{0}=1$ or not (as can be easily verified). Also we define
$b_{\tau}:=\alpha_{t}\kappa_{t},$
whenever $t=\nu^{-1}(\tau)$ for some $\tau$. With these definitions, (4.1) is equivalent to
$x_{\tau+1}=(1-b_{\tau})x_{\tau}+b_{\tau}\zeta_{\tau+1},\;\forall\tau\geq 1,$ (4.7)
Note that, in (4.7), $b_{\tau}$ is a random variable for all $\tau\geq 1$, and that there is no $b_{0}$. To analyze the behavior of (4.7), we introduce some preliminary concepts. Let ${\mathcal{F}}_{t}$ be the $\sigma$-algebra generated by $w_{0},\kappa_{0}^{t},\xi_{1}^{t}$. With the change in time indices, define $\{{\mathcal{G}}_{\tau}\}$, where ${\mathcal{G}}_{\tau}={\mathcal{F}}_{\nu^{-1}(\tau)}$, whenever $t=\nu^{-1}(\tau)$ for some $\tau$. Then it is easy to see that $\{{\mathcal{G}}_{\tau}\}$ is also a filtration, and that
$E(x_{\tau}|{\mathcal{G}}_{\tau})=E_{t}(w_{t}|{\mathcal{F}}_{t})$
whenever $t=\nu^{-1}(\tau)$ for some $\tau$. Hence we can mimic the earlier notation and denote $E(X|{\mathcal{G}}_{\tau})$ by $E_{\tau}(X)$. Also, if (as is often the case, including the applications studied here), the original step size $\alpha_{t}\in{\mathcal{M}}({\mathcal{F}}_{t})$, then $b_{\tau}\in{\mathcal{M}}({\mathcal{F}}_{t})={\mathcal{M}}({\mathcal{F}}_{\nu^{-1}(\tau)})={\mathcal{M}}({\mathcal{G}}_{\tau})$.
Now we present a general convergence result for (4.7). Observe that $\{w_{t}\}$ is a “piecewise-constant version” of $\{x_{\tau}\}$. Hence if some conclusions are established for the $v$-process, they are also established for the $w$-process, after adjusting for the time change from $t$ to $\tau$.
Theorem 4.1.
Consider the recursion (4.7). Suppose there exist constants $\mu_{t},M_{t}$ such that
$|E_{t}(\xi_{t+1})|\leq\mu_{t}(1+|w_{t}|)\;\forall t\geq 0,$ (4.8)
$CV_{t}(\xi_{t+1})\leq M_{t}^{2}(1+w_{t}^{2}),\;\forall t\geq 0.$ (4.9)
Define
$f_{\tau}=b_{\tau}^{2}(1+2\mu_{\nu^{-1}(\tau)}^{2}+M^{2}_{\nu^{-1}(\tau)})+3b_{\tau}\mu_{\nu^{-1}(\tau)},$ (4.10)
$g_{\tau}=b_{\tau}^{2}(2\mu_{\nu^{-1}(\tau)}^{2}+M^{2}_{\nu^{-1}(\tau)})+b_{\tau}\mu_{\nu^{-1}(\tau)}.$ (4.11)
Then we have the following conclusions:
  1. 1.
    If
    $\sum_{\tau=1}^{\infty}f_{\tau}<\infty,\sum_{\tau=1}^{\infty}g_{\tau}<\infty,$ (4.12)
    then $x_{\tau}$ is bounded almost surely.
  2. 2.
    If, in addition to (4.12), we also have
    $\sum_{\tau=1}^{\infty}b_{\tau}=\infty,$ (4.13)
    then $x_{\tau}\rightarrow 0$ as $\tau\rightarrow\infty$.
  3. 3.
    If both (4.12) and (4.13) hold, then $x_{\tau}=o(\tau^{-\lambda})$ for every $\lambda<1$ such that
    $\sum_{\tau=1}^{\infty}(\tau+1)^{\lambda}g_{\tau}<\infty,$ (4.14)
    $\sum_{\tau=1}^{\infty}[b_{\tau}-\lambda\tau^{-1}]=\infty,$ (4.15)
    and in addition, there exists a $T<\infty$ such that
    $b_{\tau}-\lambda\tau^{-1}\geq 0\;\forall\tau\geq T.$ (4.16)
Proof.
The proof consists of reformulating the bounds on the error ${\boldsymbol{\xi}}_{t+1}$ in such a way that Theorems 2.5 and 2.7 apply. By assumption, we have that
$|E_{t}(\xi_{t+1})|\leq\mu_{t}(1+|w_{t}|)\;\forall t.$
In particular, when $t=\nu^{-1}(\tau)$, we have that
$|E_{\tau}(\zeta_{\tau+1})|=|E_{t}(\xi_{t+1})|\leq\mu_{t}(1+|w_{t}|)=\mu_{\nu^{-1}(\tau)}(1+|x_{\tau}|).$
It follows in an entirely analogous manner that
$CV_{\tau}(\zeta_{t+1})\leq M_{\nu^{-1}(\tau)}(1+x_{\tau}^{2}).$
With these observations, we see that Theorems 2.5 and 2.7 apply to (4.7), with the only changes being that (i) the stochastic process is scalar-valued and not vector-valued, (ii) the time index is denoted by $\tau$ and not $t$, and (iii) $\mu_{t},M_{t}$ are replaced by $\mu_{\nu^{-1}(\tau)},M_{\nu^{-1}(\tau)}$ respectively. Now the conclusions of the theorem follow from Theorems 2.5 and 2.7. ∎
Now, for the convenience of the reader, we reprise the two commonly used approaches for choosing the step size, known as a “global clock” and a “local clock” respectively. This distinction was apparently first introduced in [4]. In each case, there is a deterministic sequence $\{\beta_{t}\}_{t\geq 0}$ of step sizes. If a global clock is used, then $\alpha_{t}=\beta_{t}$ at each update, so that $b_{\tau}=\beta_{\nu^{-1}(\tau)}$. If a local clock is used, then $\alpha_{t}=\beta_{\nu_{t}}$, so that then $b_{\tau}=\beta_{\tau-1}$ The extra $-1$ in the subscript is to ensure consistency in notation. To illustrate, suppose $\kappa_{t}=1$ for all $t$. Then $\nu_{t}=t+1$, and $\nu^{-1}(\tau)=\tau-1$.
Now we begin our analysis of (4.7) with the two types of clocks. Now that Theorem 4.1 is established, the challenge is to determine when (4.13) through (4.16) (as appropriate) hold for the two choices of step sizes, namely global vs. local clocks.
Towards this end, we introduce a few assumptions regarding the update process.
  1. (U1)
    $\nu_{t}\rightarrow\infty$ as $t\rightarrow\infty$ almost surely.
  2. (U2)
    We have that
[…]
  3. (U3)
    For each sample path, there exists a constant $r>0$ such that
[…]
Observe that all three assumptions are sample-pathwise. Thus (U3) implies (U2) which in turn implies (U1).
We begin by stating the convergence results when a local clock is used.
Theorem 4.2.
Suppose a local clock is used, so that $\alpha_{t}=\beta_{\nu_{t}}$, so that $b_{\tau}=\beta_{\tau-1}$. Suppose further that Assumption (U1) holds, and moreover
  1. (a)
    $\{\mu_{t}\}$ is nonincreasing; that is, $\mu_{t+1}\leq\mu_{t},\;\forall t$.
  2. (b)
    $M_{t}$ is uniformly bounded, say by $M$.
With these assumptions,
  1. 1.
    If
    $\sum_{t=0}^{\infty}\beta_{t}^{2}<\infty,\sum_{t=0}^{\infty}\beta_{t}\mu_{t}<\infty,$ (4.19)
    then $\{x_{\tau}\}$ is bounded almost surely, and $\{w_{t}\}$ is bounded almost surely.
  2. 2.
    If, in addition
    $\sum_{t=0}^{\infty}\beta_{t}=\infty,$ (4.20)
    then $x_{\tau}\rightarrow 0$ as $t\rightarrow\infty$ almost surely, and $w_{t}\rightarrow 0$ as $t\rightarrow\infty$ almost surely.
  3. 3.
    Suppose $\beta_{t}=O(t^{-(1-\phi)})$, for some $\phi>0$, and $\beta_{t}=\Omega(t^{-(1-C)})$ for some $C\in(0,\phi]$. Suppose that $\mu_{t}=O(t^{-\epsilon})$ for some $\epsilon>0$. Then $x_{\tau}\rightarrow 0$ as $\tau\rightarrow\infty$, and $w_{t}\rightarrow 0$ as $t\rightarrow\infty$, for all $\phi<\min\{0.5,\epsilon\}$. Further, $x_{\tau}=o(\tau^{-\lambda})$, and $w_{t}=o((\nu_{t})^{-\lambda})$ for all $\lambda<\epsilon-\phi$. In particular, if $\mu_{t}=0$ for all $t$, then $x_{\tau}=o(\tau^{-\lambda})$, and $w_{t}=o((\nu_{t})^{-\lambda})$ for all $\lambda<1$.
  4. 4.
    If Assumption (U2) holds instead of (U1), then in the previous item, $w_{t}=o((\nu_{t})^{-\lambda})$ can be replaced by $w_{t}=o(t^{-\lambda})$.
Proof.
The proof consists of showing that, under the stated hypotheses, the appropriate conditions in (4.12) through (4.16) hold.
Recall that $b_{\tau}=\beta_{\tau-1}$. Also, by Assumption (U1), $\nu_{t}\rightarrow\infty$ as $t\rightarrow\infty$, almost surely. Hence $\nu^{-1}(\tau)$ is well-defined for all $\tau\geq 1$.
Henceforth all arguments are along a particular sample path, and we omit the phrase “almost surely,” and also do not display the argument $\omega\in\Omega$.
We first prove Item 1 of the theorem. Recall the definitions of $f_{\tau}$ and $g_{\tau}$ from (4.10) and (4.11) respectively. Item 1 is established if (4.12) holds. For this purpose, note that $\mu_{s}\leq\mu_{t}$ if $s>t$, and $M_{t}\leq M$ for all $t$. We analyze each of the three terms comprising $f_{\tau}$. First,
$\sum_{\tau=1}^{\infty}b_{\tau}^{2}=\sum_{\tau=1}^{\infty}\beta_{\tau-1}^{2}=\sum_{t=0}^{\infty}\beta_{t}^{2}<\infty.$
Next, since $M_{t}\leq M$ for all $t$, we have that
$\sum_{\tau=1}^{\infty}b_{\tau}^{2}M_{\nu^{-1}(\tau)}^{2}\leq M^{2}\sum_{\tau=1}^{\infty}b_{\tau}^{2}<\infty.$
Finally,
$\sum_{\tau=1}^{\infty}b_{\tau}\mu_{\nu^{-1}(\tau)}\leq\sum_{\tau=1}^{\infty}\beta_{\tau-1}\mu_{\tau-1}=\sum_{t=0}^{\infty}\beta_{t}\mu_{t}<\infty.$
Here we use the fact that $\nu^{-1}(\tau)\geq\tau-1$, so that $\mu_{\nu^{-1}(\tau)}\leq\mu_{\tau-1}$. Thus (4.10) is established. Next, since $\{b_{\tau}\mu_{\nu^{-1}(\tau)}\}\in\ell_{1}$, so is $\{b_{\tau}^{2}\mu_{\nu^{-1}(\tau)}^{2}\}$. Hence (4.11) holds. This establishes that $\{x_{\tau}\}$ is bounded, which in turn implies that $\{w_{t}\}$ is bounded.
To prove Item 2, note that
$\sum_{\tau=1}^{\infty}b_{\tau}=\sum_{\tau=0}^{\infty}\beta_{\tau}=\infty.$
Hence (4.12) holds, and $x_{\tau}\rightarrow 0$ as $\tau\rightarrow\infty$, which in turn implies that $w_{t}\rightarrow 0$ as $t\rightarrow\infty$.
Finally we come to the rates of convergence. Recall that $\mu_{t}=O(t^{-\epsilon})$ while $M_{t}$ is bounded by $M$. Also, $\beta_{t}$ is chosen to be $O(t^{-(1-\phi)})$ and $\Omega(t^{-(1-C)})$. From the above, it is clear that
$f_{\tau}=O(\tau^{-2+2\phi})+O(\tau^{-1+\phi-\epsilon}).$
Hence (4.12) holds if
$-2+2\phi<-1\mbox{ and }-1+\phi-\epsilon<-1,\mbox{ or }\phi<\min\{0.5,\epsilon\}.$
Next, from the definition of $g_{\tau}$ in (4.11), it follows that
$(\nu^{-1}(\tau)+1))^{\lambda}g_{\tau}\leq(\nu^{-1}(\tau+1))^{\lambda}g_{\tau}=O(\tau^{-1+\phi-\epsilon+\lambda}).$
Hence (4.14) holds if
$-1+\phi-\epsilon+\lambda<-1\;\Longrightarrow\;\lambda<\epsilon-\phi.$
Combining everything shows that $x_{\tau}=o(\tau^{-\lambda})$ whenever
$\phi<\min\{0.5,\epsilon\},\lambda<\epsilon-\phi.$
If $\mu_{t}=0$ for all $t$, then $\epsilon$ can be chosen to be arbitrarily large. However, the limiting factor is that the argument in Theorem 2.7 holds only for $\lambda\leq 1$. Hence $x_{\tau}=o(\tau^{-\lambda})$ whenever
$\phi<0.5,\lambda<1.$
Now suppose Assumption (U2) holds, and fix some $e>0$. Then along almost all sample paths, for sufficiently large $T$ we have that $\nu_{t}/t\geq r-\epsilon$ for all $t\geq T$. Thus, whenever $t\geq T$, we have that
$\nu_{t}\geq rt\;\Longrightarrow\;o((\nu_{t})^{-\lambda})\leq o((rt)^{-\lambda})=o(t^{-\lambda}).$
Thus $w_{t}$ has the same rate of convergence as $x_{\tau}$. ∎
Since the analysis can commence after a finite number of iterations, it is easy to see that Assumption (a) above can be replaced by the following: $\{\mu_{t}\}$ is eventually nonincreasing; that is, there exists a $T<\infty$ such that
$\mu_{t+1}\leq\mu_{t},\;\forall t\geq T.$
Theorem 4.3.
Suppose a global clock is used, so that $\alpha_{t}=\beta_{t}$ whenever $t=\nu^{-1}(\tau)$ for some $\tau$ and as a result $b_{\tau}=\beta_{\nu^{-1}(\tau)}$. Suppose further that Assumption (U3) holds. Finally, suppose that $\beta_{t}$ is nonincreasing, so that $\beta_{t+1}\leq\beta_{t}$ for all $t$. $\beta_{t+1}\leq\beta_{t}$, for all $t$. Under these assumptions,
  1. 1.
    If (4.19) holds, and in addition
    $\sum_{t=0}^{\infty}\beta_{t}^{2}M_{t}^{2}<\infty,$ (4.21)
    then $\{w_{t}\}$ is bounded almost surely.
  2. 2.
    If, in addition, (4.20) holds, then $w_{t}\rightarrow 0$ as $t\rightarrow\infty$ almost surely.
  3. 3.
    Suppose in addition that $\beta_{t}=O(t^{-(1-\phi)})$, for some $\phi>0$, and $\beta_{t}=\Omega(t^{-(1-C)})$ for some $C\in(0,\phi]$. Suppose that $\mu_{t}=O(t^{-\epsilon})$ for some $\epsilon>0$, and $M_{t}=O(t^{\delta})$ for some $\delta\geq 0$. Then $w_{t}\rightarrow 0$ as $t\rightarrow\infty$ whenever
    $\phi<\min\{0.5-\delta,\epsilon\}.$
    Moreover, $w_{t}=o(t^{-\lambda})$ for all $\lambda<\epsilon-\phi$. In particular, if $\mu_{t}=0$ for all $t$, then $w_{t}=o(t^{-\lambda})$ for all $\lambda<1$.
The proof of Theorem 4.3 makes use of the following auxiliary lemma.
Lemma 4.4.
Suppose the update process $\{\kappa_{t}\}$ satisfies Assumption (U3). Suppose $\{\beta_{t}\}$ is an ${\mathbb{R}}_{+}$-valued sequence of deterministic constants such that $\beta_{t+1}\leq\beta_{t}$ for all $t$, and in addition, (4.20) holds. Then
$\sum_{\tau=1}^{\infty}\beta_{\nu^{-1}(\tau)}=\sum_{t=0}^{\infty}\beta_{t}\kappa_{t}=\infty.$ (4.22)
Proof.
We begin by showing that there exists an integer $M$ such that, whenever $2^{k}>M$, we have
$\frac{1}{2^{k}}\left(\sum_{t=2^{k}+1}^{2^{k+1}}\kappa_{t}\right)\geq\frac{r}{2}.$ (4.23)
By assumption, the ratio $\nu_{t}/t\rightarrow r$ as $t\rightarrow\infty$, where $r$ could depend on the sample path (though the dependence on $\omega$ is not displayed). So we can define $\epsilon=r/2$, and choose an integer $M$ such that
$\left|\frac{1}{T}\sum_{t=0}^{T-1}\kappa_{t}-r\right|=\left|\frac{1}{T}\sum_{t=0}^{T-1}(\kappa_{t}-r)\right|<\frac{\epsilon}{3},\;\forall T\geq M.$
Thus, if $2^{k}>M$, we have that
$\left|\frac{1}{2^{k}}\sum_{t=2^{k}+1}^{2^{k+1}}(\kappa_{t}-r)\right|$$\leq$ $\left|\frac{1}{2^{k}}\sum_{t=1}^{2^{k+1}}(\kappa_{t}-r)\right|+\left|\frac{1}{2^{k}}\sum_{t=1}^{2^{k}}(\kappa_{t}-r)\right|$
$<$ $\frac{2}{3}\epsilon+\frac{1}{3}\epsilon=\epsilon=\frac{r}{2}.$
Next, suppose that $\beta_{t+1}\leq\beta_{t}$ for all $t$. (If this holds only for all sufficiently large $t$, we just start all the summations from the time when the above holds.)
$\sum_{t=0}^{\infty}\beta_{t}\kappa_{t}$$\geq$ $\sum_{k=1}^{\infty}\left(\sum_{t=2^{k}+1}^{2^{k+1}}\beta_{t}\kappa_{t}\right)\geq\sum_{k=1}^{\infty}\left(\sum_{t=2^{k}+1}^{2^{k+1}}\beta_{2^{k+1}}\kappa_{t}\right)$
$=$ $\sum_{k=1}^{\infty}\beta_{2^{k+1}}\left(\sum_{t=2^{k}+1}^{2^{k+1}}\kappa_{t}\right)\geq\sum_{k=1}^{\infty}\beta_{2^{k+1}}2^{k}\frac{r}{2}=\frac{r}{4}\sum_{k=1}^{\infty}\beta_{2^{k+1}}2^{k+1}$
$=$ $\frac{r}{4}\sum_{k=1}^{\infty}\sum_{t=2^{k+1}+1}^{2^{k+2}}\beta_{2^{k+1}}\geq\frac{r}{4}\sum_{k=1}^{\infty}\sum_{t=2^{k+1}+1}^{2^{k+2}}\beta_{t}=\frac{r}{4}\sum_{k=5}^{\infty}\beta_{t}=\infty.$
This is the desired conclusion. ∎
Proof.
Of Theorem 4.3: Recall that a global clock is used, so that $b_{\tau}=\beta_{\nu^{-1}(\tau)}$. Hence
$\sum_{\tau=1}^{\infty}f_{\tau}$ $=$ $\sum_{\tau=1}^{\infty}[\beta_{\nu^{-1}(\tau)}^{2}+\beta_{\nu^{-1}(\tau)}^{2}M_{\nu^{-1}(\tau)}^{2}+\beta_{\nu^{-1}(\tau)}\mu_{\nu^{-1}(\tau)}]$
$=$ $\sum_{t=0}^{\infty}[\beta_{t}^{2}+\beta_{t}M_{t}^{2}+\beta_{t}\mu_{t}]<\infty$
Via entirely similar reasoning, it follows that $\{g_{\tau}\}\in\ell_{1}$. Hence (4.12) holds, and Item 1 follows.
To prove Item 2, it is necessary to establish (4.13), which in this case becomes
$\sum_{\tau=1}^{\infty}\beta_{\nu^{-1}(\tau)}=\infty.$
This is (4.13). Hence Item 2 follows.
Finally we come to the rates of convergence. The only difference is that now $M_{t}=O(t^{\delta})$ whereas it was bounded in Theorem 4.2. To avoid tedious repetition, we indicate only the changed steps. The only change is that now
$f_{\tau}=O(\tau^{-2+2\phi})+O(\tau^{-2+2\phi+2\delta})+O(\tau^{-1+\phi-\epsilon}).$
Hence (4.12) holds if
$-2+2\phi<-1,-2+2\phi+2\delta<-1,\mbox{ and }-1+\phi-\epsilon<-1,$
or
$\phi<\min\{0.5-\delta,\epsilon\}.$
Next, from the definition of $g_{\tau}$ in (4.11), it follows that
$(\nu^{-1}(\tau)+1))^{\lambda}g_{\tau}\leq(\nu^{-1}(\tau+1))^{\lambda}g_{\tau}=O(\tau^{-1+\phi-\epsilon+\lambda}).$
Hence (4.14) holds if
$-1+\phi-\epsilon+\lambda<-1\;\Longrightarrow\;\lambda<\epsilon-\phi.$
Hence $x_{\tau}=o(\tau^{-\lambda})$ and $w_{t}=o(t^{-\lambda})$ whenever
$\phi<\min\{0.5-\delta,\epsilon\},\lambda<\epsilon-\phi.$
If $\mu_{t}=0$ for all $t$, then we can choose $\epsilon$ to be arbitrarily large, and we are left with
$\phi<0.5-\delta,\lambda<1.$
4.2 Boundedness of Iterations
In this problem we state the precise problem to be studied in this subsection and the next, which is to prove the convergence, together with estimates for the rate of convergence, of BASA applied to find the fixed point of a contractive map. In this subsection, we establish the almost sure boundedness of the iterations under a very general set of assumptions, in which the contractive map is allowed to have memory, delays etc. In the next subsection, we prove the convergence of the iterations, under more restrictive assumptions; but these are still sufficiently general to be applicable to the problem of $Q$-Learning, to be studied in Section 5.
Let ${\mathbb{N}}$ denote the set of natural numbers including zero. There is a measurement function ${\bf h}:{\mathbb{N}}\times({\mathbb{R}}^{d})^{\mathbb{N}}\rightarrow({\mathbb{R}}^{d})^{\mathbb{N}}$, which is assumed to be nonanticipative [^6] In control and system theory, such a function is also referred to as “causal.” and to have finite memory. The nonanticipativeness of ${\bf h}$ means that
${\boldsymbol{\theta}}_{0}^{\infty},{\boldsymbol{\phi}}_{0}^{\infty}\in({\mathbb{R}}^{d})^{\mathbb{N}},{\boldsymbol{\theta}}_{0}^{t}={\boldsymbol{\phi}}_{0}^{t}\;\Longrightarrow\;{\bf h}(\tau,{\boldsymbol{\theta}}_{0}^{\infty})={\bf h}(\tau,{\boldsymbol{\phi}}_{0}^{\infty}),0\leq\tau\leq t.$ (4.24)
In other words, ${\bf h}(t,{\boldsymbol{\theta}}_{0}^{\infty})$ depends only on ${\boldsymbol{\theta}}_{0}^{t}$. The finite memory of ${\bf h}$ means that there exists a finite constant $\Delta$ which does not depend on $t$, such that ${\bf h}(t,{\boldsymbol{\theta}}_{0}^{t})$ further depends only on ${\boldsymbol{\theta}}_{t-\Delta+1}^{t}$. With slightly sloppy notation, this can be written as
${\bf h}(t,{\boldsymbol{\theta}}_{0}^{t})={\bf h}(t,{\boldsymbol{\theta}}^{t}_{t-\Delta+1}),\;\forall t\geq\Delta,\;\forall{\boldsymbol{\theta}}_{0}^{\infty}\in({\mathbb{R}}^{d})^{\mathbb{N}}.$ (4.25)
This formulation incorporates the possibility of “delayed information” of the form
$\eta_{t,i}=g_{i}(\theta_{1}(t-\Delta_{1}(t)),\cdots,\theta_{d}(t-\Delta_{d}(t))),$ (4.26)
where $\Delta_{1}(t),\cdots,\Delta_{d}(t)$ are delays that could depend on $t$. The only requirement is that each $\Delta_{j}(t)\leq\Delta$ for some finite $\Delta$. This formulation is analogous to [38, Eq. (2)] and [4, Eq. (1.4)], which is slightly more general in that they require only that $t-\Delta_{i}(t)\rightarrow\infty$ as $t\rightarrow\infty$, for each index $i\in[d]$. In particular, if ${\bf h}$ is “memoryless” in the sense that, for some function ${\bf g}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$, we have
${\bf h}(t,{\boldsymbol{\theta}}_{0}^{t})={\bf g}({\boldsymbol{\theta}}_{t}),$ (4.27)
then we can take $\Delta=1$.
To proceed further, it is assumed that the measurement function satisfies the following assumption:
  1. (F1)
    There exist an integer $\Delta\geq 1$ and a constant $\gamma\in(0,1)$ such that
    $\|{\bf h}(t,{\boldsymbol{\psi}}_{t-\Delta+1}^{t})-{\bf h}(t,{\boldsymbol{\phi}}_{t-\Delta+1}^{t})\|_{\infty}\leq\gamma\|{\boldsymbol{\psi}}_{t-\Delta+1}^{t}-{\boldsymbol{\phi}}_{t-\Delta+1}^{t}\|_{\infty},\;\forall t\geq\Delta,\;\forall{\boldsymbol{\psi}}_{0}^{\infty},{\boldsymbol{\phi}}_{0}^{\infty}\in({\mathbb{R}}^{d})^{\mathbb{N}}.$ (4.28)
    This assumption means that the map ${\boldsymbol{\theta}}^{t}_{t-\Delta+1}\mapsto{\bf h}(t,{\boldsymbol{\theta}}^{t}_{t-\Delta+1})$ is a contraction with respect to $\|\cdot\|_{\infty}$. In case $\Delta=1$ and ${\bf h}$ is of the form (4.27), Assumption (F1) says that the map ${\bf g}$ is a contraction.
Now we discuss a few implications of Assumption (F1).
  1. (F2)
    By repeatedly applying (4.28) over blocks of width $\Delta$, one can conclude that
    $\|{\bf h}(t,{\boldsymbol{\psi}}_{t-\Delta+1}^{t})-{\bf h}(t,{\boldsymbol{\phi}}_{t-\Delta+1}^{t})\|_{\infty}\leq\gamma^{\lfloor t/\Delta\rfloor}\|{\boldsymbol{\psi}}_{0}^{\Delta-1}-{\boldsymbol{\phi}}_{0}^{\Delta-1}\|_{\infty},\;\forall{\boldsymbol{\psi}}_{0}^{\infty},{\boldsymbol{\phi}}_{0}^{\infty}\in({\mathbb{R}}^{d})^{\mathbb{N}}.$ (4.29)
    Therefore, for every sequence ${\boldsymbol{\phi}}_{0}^{\infty}$, the iterations ${\bf h}(t,{\boldsymbol{\phi}}_{0}^{t})$ converge to a unique fixed point ${\boldsymbol{\pi}}^{*}$. In particular, if we let $({\boldsymbol{\pi}}^{*})_{0}^{\infty}$ denote the sequence whose value is ${\boldsymbol{\pi}}^{*}$ for every $t$, then it follows that
    $\|{\bf h}(t,({\boldsymbol{\pi}}^{*})_{0}^{t})-{\boldsymbol{\pi}}^{*}\|_{\infty}\leq C_{0}\gamma^{\lfloor t/\Delta\rfloor},\;\forall t,$ (4.30)
    for some constant $C_{0}$.
  2. (F3)
    The following also follows from Assumption (F1): There exist constants $\rho<1$ and $c_{1}^{\prime}>0$ such that
    $\|{\bf h}(t,{\boldsymbol{\phi}}_{0}^{t})\|_{\infty}\leq\rho\max\{c_{1}^{\prime},\|{\boldsymbol{\phi}}_{0}^{t}\|_{\infty}\},\;\forall{\boldsymbol{\phi}}\in({\mathbb{R}}^{d})^{\mathbb{N}},t\geq 0.$ (4.31)
Next, for convenience, we define
${\boldsymbol{\eta}}_{t}={\bf h}(t,{\boldsymbol{\theta}}_{0}^{t}).$ (4.32)
Strictly speaking we should write ${\boldsymbol{\eta}}_{t}={\bf h}({\boldsymbol{\theta}}_{t-\Delta+1}^{t})$, but the above notation is simpler. In order to determine ${\boldsymbol{\pi}}^{*}$ in (F2), we use BASA. Specifically, we choose ${\boldsymbol{\theta}}_{0}$ as we wish (either deterministically or at random). At time $t$, we update ${\boldsymbol{\theta}}_{t}$ to ${\boldsymbol{\theta}}_{t+1}$ according to
${\boldsymbol{\theta}}_{t+1}={\boldsymbol{\theta}}_{t}+{\boldsymbol{\alpha}}_{t}\circ[{\boldsymbol{\eta}}_{t}+{\boldsymbol{\xi}}_{t+1}],$ (4.33)
where ${\boldsymbol{\alpha}}_{t}$ is the vector of step sizes belonging to $[0,1)^{d}$, ${\boldsymbol{\xi}}_{t+1}$ is the measurement noise vector belonging to ${\mathbb{R}}^{d}$, and $\circ$ denotes the Hadamard product. We are interested in studying two questions:
  1. (Q1)
    Under what conditions is the sequence of iterations $\{{\boldsymbol{\theta}}_{t}\}$ bounded almost surely?
  2. (Q2)
    Under what conditions does the sequence of iterations $\{{\boldsymbol{\theta}}_{t}\}$ converge to ${\boldsymbol{\pi}}^{*}$ as $t\rightarrow\infty$?
In order to study the above two questions, we make some assumptions about various entities in (4.33). Let ${\mathcal{F}}_{t}$ denote the $\sigma$-algebra generated by the random variables ${\boldsymbol{\theta}}_{0}$, ${\boldsymbol{\xi}}_{1}^{t}$, and $\alpha_{0,i}^{t,i}$ for $i\in[d]$. Then it is clear that $\{{\mathcal{F}}_{t}\}$ is a filtration. As before, we denote $E(X|{\mathcal{F}}_{t})$ by $E_{t}(X)$.
The first set of assumptions in on the noise.
  1. (N1)
    There exists a finite constant $c_{1}^{\prime}$ and a sequence of constants $\{\mu_{t}\}$ such that
    $\|E_{t}({\boldsymbol{\xi}}_{t+1})\|_{2}\leq c_{1}^{\prime}\mu_{t}(1+\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty}),\;\forall t\geq 0.$ (4.34)
  2. (N2)
    There exists a finite constant $c_{2}^{\prime}$ and a sequence of constants $\{M_{t}\}$ such that
    $CV_{t}({\boldsymbol{\xi}}_{t+1})\leq c_{2}^{\prime}M_{t}^{2}(1+\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty}^{2}),\;\forall t\geq 0,$ (4.35)
    where, as before,
    $CV_{t}({\boldsymbol{\xi}}_{t+1})=E_{t}(\|{\boldsymbol{\xi}}_{t+1}-E_{t}({\boldsymbol{\xi}}_{t+1})\|_{2}^{2})$
Before proceeding further, let us compare the conditions (4.34) and (4.35) with their counterparts (2.14) and (2.15) in Theorem 2.5. It can be seen that the above two requirements are more liberal (i.e., less restrictive) than in Theorem 2.5, because the quantity $\|{\boldsymbol{\theta}}_{t}\|_{2}$ is replaced by $\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty}$. Hence, in (4.34) and (4.35), the bounds are more loose. However, Theorems 4.5 and 4.10 in the next section apply only to contractive mappings. Hence Theorems 4.5 and 4.10 complement Theorem 2.5, and do not subsume it.
The next set of assumptions is on the step size sequence.
  1. (S1)
    The random step size sequences $\{\alpha_{t,i}\}$ and the sequences $\{\mu_{t}\}$, $\{M^{2}_{t}\}$ and satisfy (almost surely)
    $\sum_{t=0}^{\infty}\alpha_{t,i}^{2}<\infty,\sum_{t=0}^{\infty}M_{t}^{2}\alpha_{t,i}^{2}<\infty,\sum_{t=0}^{\infty}\mu_{t}\alpha_{t,i}<\infty,\;\forall i\in[d].$ (4.36)
  2. (S2)
    The random step size sequence $\{\alpha_{t,i}\}$ satisfies (almost surely)
    $\sum_{t=0}^{\infty}\alpha_{t,i}=\infty,\mbox{ a.s.},\;\forall i\in[d].$ (4.37)
With these assumptions in place, we state the main result of this subsection, namely, the almost sure boundedness of the iterations. In the next subsectioon, we state and prove the convergence of the iterations. Actually, we could combine both statements into one mega-theorem. But the proof would be too long; hence both the theorem statement and the proof are split over two subsections.
Theorem 4.5.
Suppose that Assumptions (N1) and (N2) about the noise sequence, (S1) and (S2) about the step size sequence, and (F1) about the function ${\bf h}$ hold, and that ${\boldsymbol{\theta}}_{t+1}$ is defined via (4.33). Then $\sup_{t}\|{\boldsymbol{\theta}}_{t}\|_{\infty}<\infty$ almost surely.
The proof of the theorem is fairly long and involves several preliminary results and observations. Throughout, the discussion, the reader should keep in mind that Assumptions (N1), (N2), (F1) and (S1) are assumed to hold (but not (S2)).
To aid in proving the results, we introduce a sequence of “renormalizing constants.” This is similar to the technique used in [38]. For $t\geq 0$, define
$\Lambda_{t}:=\max\{\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty},c_{1}^{\prime}\},$ (4.38)
where $c_{1}^{\prime}$ is defined in (4.32). With this definition, it follows from (4.32) that ${\boldsymbol{\eta}}_{t}={\bf h}(t,{\boldsymbol{\theta}}_{0}^{t})$ satisfies
$\|{\boldsymbol{\eta}}_{t}\|_{\infty}\leq\rho\Lambda_{t},\;\forall t.$ (4.39)
Define ${\boldsymbol{\zeta}}_{t+1}=\Lambda_{t}^{-1}{\boldsymbol{\xi}}_{t+1}$ for all $t\geq 0$. Now observe that $\Lambda_{t}^{-1}\leq c_{1}^{-1}$, and $\Lambda_{t}^{-1}\leq(\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty})^{-1}$. Hence
$\|E_{t}(\zeta_{t+1,i})\|_{\infty}\leq c_{1}^{\prime}\mu_{t}(c_{1}^{-1}+1)=:c_{2}\mu_{t},$ (4.40)
where $c_{2}=c_{1}^{\prime}(c_{1}^{-1}+1)$. In particular, the above implies that
$|E_{t}(\zeta_{t+1,i})|\leq c_{2}\mu_{t},\;\forall t\geq 0.$ (4.41)
Similarly
$CV_{t}(\zeta_{t+1,i})\leq c_{3}M_{t}^{2},\;\forall t\geq 0,$ (4.42)
for some constant $c_{3}$.
If we compare (4.40) with (4.34), and (4.41) with (4.35), we see that the bounds for the “modified” error ${\boldsymbol{\zeta}}_{t+1}$ are simpler than those for ${\boldsymbol{\xi}}_{t+1}$. Specifically, the right side of both (4.40) and (4.41) are bounded with respect to ${\boldsymbol{\theta}}_{0}^{t}$ for each $t$, though they may be unbounded as functions of $t$. In contrast, the right sides of (4.34) an (4.35) are permitted to be functions of $\|{\boldsymbol{\theta}}_{0}^{t}\|_{\infty}$.
Though the next result is quite obvious, we state it separately, because it is used repeatedly in the sequel.
Lemma 4.6.
For $i\in[d]$ and $0\leq s\leq k<\infty$, define the doubly-indexed stochastic process
$D_{i}(s,k+1)=\sum_{t=s}^{k}\Bigl{[}\prod_{r=t+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{t,i}\zeta_{t+1,i},$ (4.43)
where an empty product is taken as $1$. Then $\{D_{i}(s,k)\}$ satisfies the recursion
$D_{i}(s,k+1)=(1-\alpha_{k,i})D_{i}(s,k)+\alpha_{k,i}\zeta_{k+1,i},D_{i}(s,s)=0.$ (4.44)
In the other direction, (4.43) gives a closed-form solution for the recursion (4.44).
[…]
Lemma 4.7.
There exists $\Omega_{1}\subset\Omega$ with $P(\Omega_{1})=1$ and $r_{1}^{*}:\Omega_{1}\times(0,1)\rightarrow{\mathbb{N}}$ such that
$|D_{i}(s,k+1)(\omega)|\leq\epsilon,\;\forall s\leq k,i\in[d],k\geq r_{1}^{*}(\omega,\epsilon),\omega\in\Omega_{1}.$ (4.45)
Proof.
Let $\epsilon>0$ be given. It follows from Lemma 4.6 that $D_{i}$ satisfies the recursion
$D_{i}(0,t+1)=(1-\alpha_{t,i})D_{i}(0,t)+\alpha_{t,i}\zeta_{t+1,i}$
with $D_{i}(0,0)=0$. Let us fix an index $i\in[d]$, and invoke (4.41) and (4.42). Then it follows from (4.42) that
$CV_{t}(\zeta_{t+1,i})\leq c_{3}M_{t}^{2},$
and (4.41) also holds. Now, if Assumptions (S1) and (S2) also hold, then all the hypotheses needed to apply Theorem 2.5 are in place. Therefore $D_{i}(0,k+1)$ converges to zero almost surely. This holds for each $i\in[d]$ Therefore, if we define
$\Omega_{1}=\{\omega\in\Omega_{1}:D_{i}(0,k+1)(\omega)\rightarrow 0\mbox{ as }t\rightarrow\infty\;\forall i\in[d]\},$
then $P(\Omega_{1})=1.$ We can see that for $\omega\in\Omega_{1}$ we can choose $r_{1}^{*}(\omega,\epsilon)\text{ such that }\;\forall k\geq r_{1}^{*}(\omega,\epsilon),i\in[d]$ we have
$|D_{i}(0,k+1)(\omega)|\leq\textstyle\frac{1}{2}\epsilon.$
Note that for $s\leq k$
$D_{i}(0,k+1)(\omega)=\left[\prod_{r=s}^{k}(1-\alpha_{r,i}(\omega))\right]D_{i}(0,s)(\omega)+D_{i}(s,k+1)(\omega).$
As a consequence
$|D_{i}(s,k+1)(\omega)|\leq|D_{i}(0,k+1)(\omega)|+|D_{i}(0,s)(\omega)|.$
Thus (4.45) holds for the choice of $r_{1}^{*}(\omega,\epsilon)$ made above. ∎
Lemma 4.8.
There exists $\Omega_{2}\subset\Omega$ with $P(\Omega_{2})=1$ and $r_{2}^{*}:\Omega_{1}\times{\mathbb{N}}\times(0,1)\rightarrow{\mathbb{N}}$ such that
$\prod_{s=j}^{k}(1-\alpha_{s,i}(\omega))\leq\epsilon,\;\forall\,k\geq r_{2}^{*}(\omega,j,\epsilon),i\in[d],\omega\in\Omega_{2}.$ (4.46)
Proof.
In view of the assumption (S2), if we define
$\Omega_{2}=\left\{\omega\in\Omega:\sum_{s=j}^{\infty}\alpha_{t,i}(\omega)=\infty\;\forall i\in[d]\right\},$
then $P(\Omega_{2})=1$. For all $\omega\in\Omega_{2}$, we have
$\sum_{s=j}^{\infty}\alpha_{t,i}(\omega)=\infty.$
Using the elementary inequality $(1-x)\leq\exp\{-x\}$ for all $x\in[0,\infty)$, it follows that
$\prod_{s=j}^{k}(1-\alpha_{t,i}(\omega))\leq\exp\left\{-\sum_{s=j}^{k}\alpha_{t,i}(\omega)\right\}.$
Hence for $\omega\in\Omega_{2}$, $\prod_{s=j}^{k}(1-\alpha_{t,i}(\omega))$ converges to zero as $k\rightarrow\infty$. Thus we can choose $r_{2}^{*}(\omega,j,\epsilon)$ with the required property. ∎
In the rest of this section, we will fix $\omega\in\Omega_{1}\cap\Omega_{2}$, the functions $r_{1}^{*}$, $r_{2}^{*}$ obtained in Lemma 4.7 and Lemma 4.8 respectively and prove that if (F1) holds, then $\|{\boldsymbol{\theta}}_{t}(\omega)\|_{\infty}$ is bounded, which proves Theorem 4.1.
Let us rewrite the updating rule (4.33) as
$\theta_{t+1,i}=(1-\alpha_{t,i})\theta_{t,i}+\alpha_{t,i}(\eta_{t,i}+\Lambda_{t}\zeta_{t+1,i}),i\in[d],\,t\geq 0,$ (4.47)
By recursively invoking (4.47) for $k\in[0,t]$, we get
$\theta_{t+1,i}=A_{t+1,i}+B_{t+1,i}+C_{t+1,i}$ (4.48)
where
$A_{t+1,i}=\Bigl{[}\prod_{k=0}^{t}(1-\alpha_{k,i})\Bigr{]}\theta_{0,i},$ (4.49)
$B_{t+1,i}=\sum_{k=0}^{t}\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}\eta_{k,i},$ (4.50)
$C_{t+1,i}=\sum_{k=0}^{t}\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}\Lambda_{k}\zeta_{k+1,i}.$ (4.51)
Lemma 4.9.
For $i\in[d]$ ,
$|C_{t+1,i}|\leq\Lambda_{t}\sup_{0\leq r\leq t}|D_{i}(r,t+1)|.$ (4.52)
Proof.
We begin by establishing an alternate expression for $C_{k,i}$, namely
$C_{t+1,i}=\Lambda_{0}D_{i}(0,t+1)+\sum_{k=1}^{t}(\Lambda_{k}-\Lambda_{k-1})D_{i}(k,t+1),$ (4.53)
where $D_{i}(\cdot,\cdot)$ is defined in (4.44). For this purpose, observe from Lemma 4.6 that $C_{t+1,i}$ satisfies
$C_{t+1,i}=\Lambda_{t}\alpha_{t,i}\zeta_{t+1,i}+(1-\alpha_{t,i})C_{t,i}=\Lambda_{t}D_{i}(t,t+1)+(1-\alpha_{t,i})C_{t,i},$ (4.54)
because $\alpha_{t,i}\zeta_{t+1,i}=D_{i}(t,t+1)$ due to (4.44) with $s=t$. The proof of (4.53) is by induction. It is evident from (4.51) that
$C_{1,i}=\Lambda_{0}\alpha_{0,1}\zeta_{1,i}=\Lambda_{0}D_{i}(0,1).$
Thus (4.53) holds when $t=0$. Now suppose by way of induction that
$C_{t,i}=\Lambda_{0}D_{i}(0,t)+\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})D_{i}(k,t).$ (4.55)
Using this assumption, and the recursion (4.54), we establish (4.53).
Substituting from (4.55) into (4.54) gives
$C_{t+1,i}=\Lambda_{t}D_{i}(t,t_{1})+\Lambda_{0}(1-\alpha_{t,i})D_{i}(0,t)+(1-\alpha_{t,i})\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})D_{i}(k,t).$ (4.56)
Now (4.44) implies that
$(1-\alpha_{t,i})D_{i}(k,t)=D_{i}(k,t+1)-\alpha_{t,i}\zeta_{t+1,i}=D_{i}(k,t+1)-D_{i}(t,t+1).$
Therefore the summation in (4.56) becomes
$\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})(1-\alpha_{t,i})D_{i}(k,t)$ $=$ $\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})D_{i}(k,t)$
-- $D_{i}(t,t+1)\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})=S_{1}+S_{2}\mbox{ say}.$
Then $S_{2}$ is just a telescoping sum and equals
$S_{2}=-\Lambda_{t-1}D_{i}(t,t+1)+\Lambda_{0}D_{i}(t,t+1).$
The second term in (4.56) equals
$\Lambda_{0}(1-\alpha_{t,i})D_{i}(0,t)=\Lambda_{0}[D_{i}(0,t+1)-\alpha_{t,i}\zeta_{t+1,i}]=\Lambda_{0}D_{i}(0,t+1)-\Lambda_{0}D_{i}(t,t+1).$
Putting everything together and observing that the term $\Lambda_{0}D_{i}(t,t+1)$ cancels out gives
$C_{t+1,i}=\Lambda_{0}D_{i}(0,t+1)+(\Lambda_{t}-\Lambda_{t-1})D_{i}(t,t+1)+\sum_{k=1}^{t-1}(\Lambda_{k}-\Lambda_{k-1})D_{i}(k,t).$
This is the same as (4.56) with $t+1$ replacing $t$. This completes the induction step and thus (4.53) holds. Using the fact that $\Lambda_{t}\geq\Lambda_{t-1}$, the desired bound (4.52) follows readily. ∎
Proof.
(Of Theorem 4.1) As per the statement of the theorem, we assume that (F1) holds. We need to prove that
$\sup_{t\geq 0}\Lambda_{t}<\infty.$
Define
$\delta=\min\{\frac{1-\rho}{2\rho},\frac{1}{2}\},$
and observe that, as a consequence, we have that $\rho(1+2\delta)\leq 1$. Choose $r_{1}^{*}=r_{1}^{*}(\delta)$ as in Lemma 4.8 such that
$|D_{i}(s,k+1)|\leq\delta\;\forall k\geq s\geq r_{1}^{*},\;\forall i\in[d].$
It is now shown that
$\Lambda_{t}\leq(1+2\delta)\Lambda_{r_{1}^{*}}\;\forall t,\;\forall i\in[d].$ (4.57)
By the monotonicity of $\{\Lambda_{t}\}$, it is already known that $\Lambda_{t}\leq\Lambda_{r_{1}^{*}}$ for $t\leq r_{1}^{*}$. Hence, once (4.57) is established, it will follow that
$\sup_{0\leq t<\infty}\Lambda_{t}\leq(1+2\delta)\Lambda_{r_{1}^{*}}.$
The proof of (4.57) is by induction on $t$. Accordingly, suppose (4.57) holds for $t\leq k$. Using (4.52), we have
$|C_{k+1,i}|\leq\delta\Lambda_{k}\leq\Lambda_{r_{1}^{*}}\delta(1+2\delta).$ (4.58)
It is easy to see from its definition that
$|A_{k+1,i}|\leq\Lambda_{r_{1}^{*}}\Bigl{[}\prod_{s=0}^{k}(1-\alpha_{s,i})\Bigr{]}$
Using the induction hypothesis that $\Lambda_{t}\leq(1+2\delta)\Lambda_{r_{1}^{*}}$ for $t\leq k$, we have
$\begin{split}|B_{k+1,i}|&\leq\sum_{s=0}^{k}\Bigl{[}\prod_{r=s+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{s,i}|\eta_{s,i}|\\ &\leq\sum_{s=0}^{k}\Bigl{[}\prod_{r=s+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{s,i}\rho\Lambda_{s}\\ &\leq\rho(1+2\delta)\Lambda_{r_{1}^{*}}\sum_{s=0}^{k}\Bigl{[}\prod_{r=s+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{s,i}\\ &\leq\Lambda_{r_{1}^{*}}\sum_{s=0}^{k}\Bigl{[}\prod_{r=s+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{s,i},\end{split}$
because $\rho(1+2\delta)\leq 1$. Also, the following identity is easy to prove by induction.
$\Bigl{[}\prod_{s=0}^{k}(1-\alpha_{s,i})\Bigr{]}+\sum_{s=0}^{k}\Bigl{[}\prod_{r=s+1}^{k}(1-\alpha_{r,i})\Bigr{]}\alpha_{s,i}=1\;\forall k<\infty$ (4.59)
Combining these bounds gives
$|A_{k+1,i}|+|B_{k+1,i}|\leq\Lambda_{r_{1}^{*}}.$
Combining this with (4.48) and (4.58) leads to
$\theta_{k+1,i}\leq\Lambda_{r_{1}^{*}}(1+\delta(1+2\delta))\leq\Lambda_{r_{1}^{*}}(1+2\delta).$
Therefore $\|{\boldsymbol{\theta}}_{k+1}\|_{\infty}\leq\Lambda_{r_{1}^{*}}(1+2\delta)$, and
$\Lambda_{k+1}=\max\{\|{\boldsymbol{\theta}}_{k+1}\|_{\infty},\Lambda_{k}\}\leq\Lambda_{r_{1}^{*}}(1+2\delta).$
This proves the induction hypothesis completes the proof of Theorem 4.1. ∎
4.3 Convergence of Iterations with Rates
In this subsection, we further study the iteration sequence (4.33), under a variety of Block (or Batch) updating schemes, corresponding to various choices of the step sizes. Whereas the almost sure boundedness of the iterations is established in the previous subsection, in this subsection we prove that the iterations converge to the desired fixed point ${\boldsymbol{\pi}}^{*}$. Then we also find bounds on the rate of convergence.
We study three specific methods for choosing the step size vector ${\boldsymbol{\alpha}}_{t}$ in (4.33). Within the first two methods, we further divide into local clocks and global clocks. However, in the third method, we permit only the use of a global clock, for reasons to be specified.
4.3.1 Convergence Theorem
The overall plan is to follow up Theorem 4.5, which establishes the almost sure boundedness of the iterations, with a stronger result showing that the iterations converge almost surely to ${\boldsymbol{\pi}}^{*}$, the fixed point of the map ${\bf h}$. This convergence is established under the same assumptions as in Theorem 4.5. In particular, the step size sequence is assumed to satisfy (S1) and (S2). Having done this, we then study when (S1) and (S2) hold for each of the three methods for choosing the step sizes.
Theorem 4.10.
Suppose that Assumptions (N1) and (N2) about the noise sequence, (S1) and (S2) about the step size sequence, and (F1) about the function ${\bf h}$ hold, and that ${\boldsymbol{\theta}}_{t+1}$ is defined via (4.33). Then ${\boldsymbol{\theta}}_{t}\rightarrow{\boldsymbol{\pi}}^{*}$ as $t\rightarrow\infty$ almost surely, where ${\boldsymbol{\pi}}^{*}$ is defined in (F2).
Proof.
From (4.48), we have an expression for ${\boldsymbol{\theta}}_{t+1,i}$, where $A_{t+1,i}$, $B_{t+1,i}$ and $C_{t+1,i}$ are given by (4.49), (4.50) and (4.51) respectively. Also, by changing notation from $k$ to $t$ and $s$ to $k$ in (4.59), and multiplying both sides by $\pi^{*}_{i}$, we can write
$\pi^{*}_{i}=\Bigl{[}\prod_{k=0}^{t}(1-\alpha_{k,i})\Bigr{]}\pi^{*}_{i}+\left\{\sum_{k=0}^{t}\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}\right\}\pi^{*}_{i},\;\forall t.$
Substituting from these formulas gives
$\theta_{t+1,i}-\pi^{*}_{i}=\bar{A}_{t+1,i}+\bar{B}_{t+1,i}+C_{t+1,i},$ (4.60)
where
$\bar{A}_{t+1,i}=\prod_{k=0}^{t}(1-\alpha_{k,i})(\theta_{0,i}-\pi^{*}_{i}),$ (4.61)
$\bar{B}_{t+1,i}=\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}(\eta_{k,i}-\pi^{*}_{i}),$ (4.62)
and $C_{t+1,i}$ is as in (4.51). It is shown in turn that each of these quantities approaches zero as $t\rightarrow\infty$.
First, from Assumption (S2), it follows that [^7] We omit the phrase “almost surely” in these arguments.
$\prod_{k=0}^{t}(1-\alpha_{k,i})\rightarrow 0\mbox{ as }t\rightarrow\infty.$
Since $\theta_{0,i}-\pi^{*}_{i}$ is a constant along each sample path, $\bar{A}_{t+1,i}$ approaches zero.
Second, by combining (4.29) and (4.30) in Property (F2), it follows that
$|\eta_{t,i}-\pi^{*}_{i}|\leq\gamma^{\lfloor t/\Delta\rfloor}\|{\boldsymbol{\theta}}_{0}^{\Delta}-({\boldsymbol{\pi}}^{*})^{\Delta}\|_{\infty}\leq C_{1}\gamma^{\lfloor t/\Delta\rfloor}$
for some constant $C_{1}$ (which depends on the sample path). Thus
$\sum_{r=0}^{\infty}|\eta_{t,i}-\pi^{*}_{i}|<\infty$
along almost all sample paths. Now it follows from (4.62) that
$|\bar{B}_{t+1,i}|$$\leq$ $\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}|\eta_{k,i}-\pi^{*}_{i}|$ (4.63)
$\leq$ $\Bigl{[}\prod_{r=k+1}^{t}(1-\alpha_{r,i})\Bigr{]}\alpha_{k,i}C_{1}\gamma^{\lfloor t/\Delta\rfloor}=:L_{t+1,i}.$
Let $L_{t+1,i}$ denote the right side of this inequality. Then it follows from Lemma 4.6 that $L_{t+1,i}$ satisfies the recursion
$L_{t+1,i}=(1-\alpha_{t,i})L_{t,i}+\alpha_{t,i}C_{1}\gamma^{\lfloor t/\Delta\rfloor}.$ (4.64)
The convergence of $L_{t+1,i}$ to zero can be proved using Theorem 4.1. Since the quantity $C_{1}\gamma^{\lfloor t/\Delta\rfloor}$ is deterministic, its mean is itself and its variance is zero. So in (4.8) and (4.9), we can define
$\mu_{t}^{L}:=C_{1}\gamma^{\lfloor t/\Delta\rfloor},M_{t}^{L}:=0\;\forall t.$
We can substitute these definitions into (4.10) and (4.11), and define
$f_{\tau}^{L}=b_{\tau}^{2}(1+2\mu_{\nu^{-1}(\tau)}^{2})+3b_{\tau}\mu_{\nu^{-1}(\tau)},$ (4.65)
$g_{\tau}^{L}=b_{\tau}^{2}(2\mu_{\nu^{-1}(\tau)}^{2})+b_{\tau}\mu_{\nu^{-1}(\tau)}.$ (4.66)
Since $\alpha_{t}\in[0,1]$ and the sequence $\{\mu_{t}^{L}\}$ is summable (because $\gamma<1$), and $M_{t}^{L}\equiv 0$, (4.12) is satisfied. Also, by Assumption (S2), (4.13) is satisfied. Hence $L_{t+1,i}\rightarrow 0$ as $t\rightarrow\infty$, which in turn implies that $\bar{B}_{t+1,i}\rightarrow 0$ as $t\rightarrow\infty$.
Finally, we come to $C_{t+1,i}$. It is evident from (4.51) and Lemma 4.6 that $C_{t+1,i}$ satisfies the recursion
$C_{t+1,i}=(1-\alpha_{t,i})C_{t,i}+\alpha_{t,i}\Lambda_{t}\zeta_{t,i}.$ (4.67)
Now observe that $\Lambda_{t}$ is bounded, and the rescaled error signal $\zeta_{t+1,i}$ satisfies (4.41) and (4.42). Hence, if $\Lambda^{*}$ is a bound for $\Lambda_{t}$, then it follows from (4.41) and (4.42) that
$|E_{t}(\Lambda_{t}\zeta_{t+1,i})|\leq c_{2}\Lambda^{*}\mu_{t},\;\forall t\geq 0,CV_{t}(\Lambda_{t}\zeta_{t+1,i})\leq c_{3}\Lambda^{*}M_{t}^{2},\;\forall t\geq 0,$ (4.68)
Hence, when Assumptions (S1) and (S2) hold, it follows from Theorem 4.1 that $C_{t+1,i}\rightarrow 0$ as $t\rightarrow\infty$. ∎
4.3.2 Various Types of Updating and Rates of Convergence
Next, we describe three different ways of choosing the update processes $\{\kappa_{t,i}\}$.
Bernoulli Updating: For each $i\in[d]$, choose a rate $\rho_{i}\in(0,1]$, and let $\{\kappa_{t,i}\}$ be a Bernoulli process such that
$\Pr\{\kappa_{t,i}=1\}=\rho_{i},\;\forall t.$
Moreover, the processes $\{\kappa_{t,i}\}$ and $\{\kappa_{t,j}\}$ are independent whenever $i\neq j$. Let $\nu_{t,i}$, the counter process for coordinate $i$, be defined as usual. Then it is easy to see that $\nu_{t,i}/t\rightarrow\rho_{i}$ as $t\rightarrow\infty$, for each $i\in[d]$. Thus Assumption (U3) is satisfied for each $i\in[d]$.
Markovian Updating: Suppose $\{Y_{t}\}$ is a sample path of an irreducible Markov process on the state space $[d]$. Define the update process $\{\kappa_{t,i}\}$ by
$\kappa_{t,i}=I_{\{Y_{t}=i\}}=\left\{\begin{array}[]{ll}1,&\mbox{if }Y_{t}=i,\\ 0,&\mbox{if }Y_{t}\neq i.\end{array}\right.$
Let ${\boldsymbol{\mu}}$ denote the stationary distribution of the Markov process. Then the ratio $\nu_{t,i}/t\rightarrow\mu_{i}$ as $t\rightarrow\infty$, for each $i\in[d]$. Hence once again Assumption (U3) holds.
Batch Markovian Updating: This is an extension of the above. Instead of a single Markovian sample path, there are $N$ different sample paths, denoted by $\{Y_{t}^{n}\}$ where $n\in[N]$. Each sample path $\{Y_{t}^{n}\}$ comes an irreducible Markov process over the state space $[d]$, and the dynamics of different Markov processes could be different (though there does not seem to be any advantage to doing this). The update process is now given by
$\kappa_{t,i}=\sum_{n\in[N]}I_{\{Y_{t}^{n}=i\}}.$
Define the counter process $\nu_{t,i}$ as before, and let ${\boldsymbol{\mu}}^{n}$ denote the stationary distribution of the $n$-th Markov process. Then
$\frac{\nu_{t,i}}{t}\rightarrow\sum_{n\in[N]}\mu_{i}^{n}.$
Hence once again Assumption (U3) holds.
Now we establish convergence rates under each of the above updating methods (and indeed, any method such that Assumption (U3) is satisfied). The proof of Theorem 4.10 gives us a hint on how this can be done. Specifically, each of the entities $\bar{A}_{t+1,i},L_{t+1,i},C_{t+1,i}$ satisfies a stochastic recursion, whose rate of convergence can be established using Theorems 4.2 and 4.3. These theorems apply to scalar-valued stochastic processes with intermittent updating. In principle, when updating ${\boldsymbol{\theta}}_{t}$, we could use a mixture of global and local clocks for different components. However, in our view, this would be quite unnatural. Instead, it is assumed that for every component, either a global clock or a local clock is used. Recall also the bounds (4.34) and (4.35) on the error ${\boldsymbol{\xi}}_{t+1}$.
Theorem 4.11.
Suppose a local clock is used, so that $\alpha_{t,i}=\beta_{\nu_{t,i}}$ for each $i$ that is updated at time $t$. Suppose that $\{\mu_{t}\}$ is nonincreasing; that is, $\mu_{t+1}\leq\mu_{t},\;\forall t$, and $M_{t}$ is uniformly bounded, say by $M$. Suppose in addition that $\beta_{t}=O(t^{-(1-\phi)})$, for some $\phi>0$, and $\beta_{t}=\Omega(t^{-(1-C)})$ for some $C\in(0,\phi]$. Suppose that $\mu_{t}=O(t^{-\epsilon})$ for some $\epsilon>0$. Then ${\boldsymbol{\theta}}_{\tau}\rightarrow 0$ as $\tau\rightarrow\infty$ for all $\phi<\min\{0.5,\epsilon\}$. Further, ${\boldsymbol{\theta}}_{\tau}=o(\tau^{-\lambda})$ for all $\lambda<\epsilon-\phi$. In particular, if $\mu_{t}=0$ for all $t$, then ${\boldsymbol{\theta}}_{\tau}=o(\tau^{-\lambda})$ for all $\lambda<1$.
The proof of the rate of convergence uses Item (3) of Theorem 4.1. In the proof, let us ignore the index $i$ wherever possible, because the subsequent analysis applies to each index $i$. Recall that $\bar{A}_{t+1,i}$ is defined in (4.61). Since $\ln(1-x)\leq-x$ for all $x\in(0,1)$, it follows that
$\ln\prod_{k=0}^{t}(1-\alpha_{k,i})\leq-\sum_{k=0}^{t}\alpha_{k,i},$
where $\alpha_{k,i}=0$ unless there is an update at time $k$. Now, since a local clock is used, we have that $\alpha_{k,i}=\beta_{\nu_{k,i}}$ whenever there is an update at time $k$. Therefore
$\sum_{k=0}^{t}\alpha_{k,i}=\sum_{s=0}^{\nu_{t,i}}\beta_{s}$
Now, if Assumption (U3) holds (which it does for each of the three types of updating considered), it follows that $\nu_{t,i}\approx t/r$ for large $t$. Thus, if $\beta_{\tau}=\Omega(\tau^{-(1-C)})$, then we can reason as follows:
$\sum_{s=0}^{\nu_{t}}\beta_{s}\approx\sum_{s=0}^{t/r}s^{-(1-C)}\approx(t/r)^{C}.$
Therefore, for large enough $t$, we have that
$\prod_{k=0}^{t}(1-\alpha_{k})\leq\exp(-(t/r)^{C}).$
It follows from (4.61) that $\bar{A}_{t+1,i}\rightarrow 0$ geometrically fast.
Next we come to $\bar{B}_{t+1,i}$, which is bounded by $L_{t+1,i}$, as defined in (4.64). Recall the definitions (4.65) and (4.66) for the sequences $\{f_{\tau}^{L}\}$ and $\{g_{\tau}^{L}\}$. Then (4.12) and (4.13) will hold whenever $C>0$. Since Assumption (U3) holds, we have that
$\mu_{\nu^{-1}(\tau)}^{L}=C_{1}\gamma^{\lfloor\nu^{-1}(\tau)/\Delta\rfloor}\leq C_{2}\gamma^{r^{\prime}\tau}$
for suitable constants $C_{2}$ and $r^{\prime}$. The point to note is that the sequence $\{C_{2}\gamma^{r^{\prime}\tau}\}$ is a geometrically convergent sequence because $\gamma<1$. Therefore (4.14) holds for every $\lambda>0$. Also, (4.15) holds for all $C>0$. Hence it follows from Item (3) of Theorem 4.1 that $L_{t+1,i}=o(t^{-\lambda})$ for every $\lambda>0$.
This leaves only $C_{t+1,i}$. We already know that $C_{t+1,i}$ satisfies the recursion (4.67). Moreover, the modified error sequence $\{\Lambda_{t}\zeta_{t,i}\}$ satisfies (4.68). The estimates for the rate of convergence now follow from Item (3) of Theorem 4.1, and need not be discussed again.
Theorem 4.12.
Suppose a global clock is used, so that $\alpha_{t,i}=\beta_{t,i}$ whenever the $i$-th component of ${\boldsymbol{\theta}}_{t}$ is updated. Suppose that $\beta_{t}$ is nonincreasing, so that $\beta_{t+1}\leq\beta_{t}$ for all $t$. Suppose in addition that $\beta_{t}=O(t^{-(1-\phi)})$, for some $\phi>0$, and $\beta_{t}=\Omega(t^{-(1-C)})$ for some $C\in(0,\phi]$. Suppose that $\mu_{t}=O(t^{-\epsilon})$ for some $\epsilon>0$, and $M_{t}=O(t^{\delta})$ for some $\delta\geq 0$. Then ${\boldsymbol{\theta}}_{t}\rightarrow 0$ as $t\rightarrow\infty$ whenever
$\phi<\min\{0.5-\delta,\epsilon\}.$
Moreover, ${\boldsymbol{\theta}}_{t}=o(t^{-\lambda})$ for all $\lambda<\epsilon-\phi$. In particular, if $\mu_{t}=0$ for all $t$, then ${\boldsymbol{\theta}}_{t}=o(t^{-\lambda})$ for all $\lambda<1$.
The proof is omitted as it is very similar to that of Theorem 4.11.
5 Applications to Reinforcement Learning
In this subsection we study the convergence of a “batch” version of the well-known $Q$-learning algorithm, first introduced in [42]. The original $Q$-learning algorithm aims to find the optimal policy for a Markov decision process (MDP), without knowing the dynamics of the underlying MDP. Instead, the learner is able to observe the sample path of the MDP, one time instant at a time, for the current state-action pair. In the original formulation, there is no guidance on how to choose the next action. There is an underlying, but often unstated, assumption that in the MDP under study, every policy must lead to an irreducible Markov process. However, as shown below, verifying this assumption is likely to be NP-hard. We propose here an alternate approaching of running $m$ different processes in parallel, where $m$ is the number of actions. In this case, the only assumption required is that each fixed action leads to an irreducible Markov process, which is easy to verify.
We begin by introducing the $Q$-learning algorithm. In the description below, we give only the essential details, and refer the reader to the standard texts [29, 36] for full details. Though this material is standard, the discussion serves to fix notation. Let ${\mathcal{X}}=\{x_{1},\cdots,x_{n}\}$ be a finite state space, and ${\mathcal{U}}=\{u_{1},\cdots,u_{m}\}$ be a finite “action space.” There are $m$ different state transition matrices $A^{u_{k}},k\in[m]$ over the state space ${\mathcal{X}}$. Thus
$a_{ij}^{u_{k}}=\Pr\{X_{t+1}=x_{j}|X_{t}=x_{i}\;\&\;U_{t}=u_{k}\}.$
In addition, there is a “reward” function $R:{\mathcal{X}}\times{\mathcal{U}}\rightarrow{\mathbb{R}}$. A policy $\pi$ is any map from ${\mathcal{X}}$ into ${\mathcal{U}}$. With each policy $\pi$, one can associate a corresponding reward function $R_{\pi}:{\mathcal{X}}\rightarrow{\mathbb{R}}$, and a Markov process with state transition matrix $A_{\pi}\in[0,1]^{n\times n}$, as follows:
$R_{\pi}(x_{i})=R(x_{i},\pi(x_{i})),(A_{\pi})_{ij}=a^{\pi(x_{i})}_{ij}.$
Given a discount factor $\gamma\in(0,1)$, we can associate an Expected Discounted Future Reward which is denoted by ${\bf v}_{\pi}\in{\mathbb{R}}^{n}$. The objective in Markov Decision Processes (MDPs) is to determine an optimal policy $\pi^{*}$ that maximizes each component of ${\bf v}_{\pi}$ over the set of all possible policies. Define
${\bf v}^{*}=\max_{\pi}{\bf v}_{\pi}.$ (5.1)
Implicit in the above definition is the fact that the same policy optimizes every single component of ${\bf v}_{\pi}$. We can also denote the $i$-th component of ${\bf v}^{*}$ by $V^{*}(x_{i})$. Then $V^{*}$ satisfies the Bellman optimality equation
$V^{*}(x_{i})=\max_{u_{k}\in{\mathcal{U}}}\left[R(x_{i},u_{k})+\gamma\sum_{j=1}^{n}a_{ij}^{u_{k}}V^{*}(x_{j})\right].$ (5.2)
Finally, any policy $\pi$ such that $V_{\pi}(x_{i})=V^{*}(x_{i})$ is an optimal policy.
The challenge thus arises as to how an optimal policy can be determined. This task is greatly simplified by the introduction of the action-value function in [42].
Theorem 5.1.
(See [42].) Define $Q^{*}:{\mathcal{X}}\times{\mathcal{U}}\rightarrow{\mathbb{R}}$ by
$Q^{*}(x_{i},u_{k})=R(x_{i},u_{k})+\gamma\sum_{j=1}^{n}a_{ij}^{u_{k}}V^{*}(x_{j}).$ (5.3)
Then $Q^{*}(\cdot,\cdot)$ satisfies the following relationships:
$Q^{*}(x_{i},u_{k})=R(x_{i},u_{k})+\gamma\sum_{j=1}^{n}a_{ij}^{u_{k}}\max_{w_{l}\in{\mathcal{U}}}Q^{*}(x_{j},w_{l}).$ (5.4)
$V^{*}(x_{i})=\max_{u_{k}\in{\mathcal{U}}}Q^{*}(x_{i},u_{k}),$ (5.5)
Moreover, every policy $\pi\in\Pi_{d}$ such that
$\pi^{*}(x_{i})=\operatornamewithlimits{arg~{}max}_{u_{k}\in{\mathcal{U}}}Q^{*}(x_{i},u_{k})$ (5.6)
is optimal.
The difficulty with the above theorem is that the characterization (5.3) is implicit, as the unknown quantity $Q^{*}$ occurs on both sides. An alternate characterization of $Q^{*}$ as the fixed point of a contraction provides a way of computing $Q^{*}$ iteratively. Define $F:{\mathbb{R}}^{|{\mathcal{X}}|\times|{\mathcal{U}}|}\rightarrow{\mathbb{R}}^{|{\mathcal{X}}|\times|{\mathcal{U}}|}$ by
$[F(Q)](x_{i},u_{k}):=R(x_{i},u_{k})+\gamma\sum_{j=1}^{n}a_{ij}^{u_{k}}\max_{w_{l}\in{\mathcal{U}}}Q(x_{j},w_{l}).$ (5.7)
Then it can be shown that the map $F$ is monotone and is a contraction with respect to the “vector $\ell_{\infty}$-norm”. Thus $Q^{*}$ is the fixed point of the map $F$. However, computing $Q^{*}$ by repeatedly applying the $F$-iteration requires full knowledge of all the parameters of the MDP. The $Q$-learning algorithm proposed in [42] does not require knowledge of the parameters of the MDP. Instead it requires only the ability to observe the next state $X_{t+1}$, once the current state $X_{t}$ is known and the current action $U_{t}$ is chosen. The algorithm begins with the characterization (5.3) of $Q^{*}$, and is an asynchronous updating procedure, as follows:
  1. 1.
    Choose an arbitrary initial guess $Q_{0}:{\mathcal{X}}\times{\mathcal{U}}\rightarrow{\mathbb{R}}$ and an initial state $X_{0}\in{\mathcal{X}}$.
  2. 2.
    At time $t$, with current state $X_{t}=x_{i}$, choose a current action $U_{t}=u_{k}\in{\mathcal{U}}$, and let the Markov process run for one time step. Observe the resulting next state $X_{t+1}=x_{j}$. Then update the function $Q_{t}$ as follows:
    $\begin{split}Q_{t+1}(x_{i},u_{k})&=Q_{t}(x_{i},u_{k})+\beta_{t}[R(x_{i},u_{k})+\gamma V_{t}(x_{j})-Q_{t}(x_{i},u_{k})],\\ Q_{t+1}(x_{s},w_{l})&=Q_{t}(x_{s},w_{l}),\;\forall(x_{s},w_{l})\neq(x_{i},u_{k}).\end{split}$ (5.8)
    where
    $V_{t}(x_{j})=\max_{w_{l}\in{\mathcal{U}}}Q_{t}(x_{j},w_{l}),$ (5.9)
    and $\{\beta_{t}\}$ is a deterministic sequence of step sizes.
  3. 3.
    Repeat.
In earlier work such as [38, 16], it is shown that the $Q$-learning algorithm converges to the optimal action-value function $Q^{*}$ provided
$\sum_{t=0}^{\infty}\beta_{t}I_{(X_{t},U_{t})=(x_{i},u_{k})}=\infty,\;\forall(x_{i},u_{k})\in{\mathcal{X}}\times{\mathcal{U}},$ (5.10)
$\sum_{t=0}^{\infty}\beta_{t}^{2}I_{(X_{t},U_{t})=(x_{i},u_{k})}<\infty,\;\forall(x_{i},u_{k})\in{\mathcal{X}}\times{\mathcal{U}}.$ (5.11)
Note that in the $Q$-learning algorithm, there is no guidance on how to choose the next action $U_{t}$. Presumably one tries to ensure that the assumptions (5.8) and (5.9) are satisfied. But that in itself does not provide a strategy for choosing $U_{t}$.
The batch $Q$-learning algorithm introduced here is as follows:
  1. 1.
    Choose an arbitrary initial guess $Q_{0}:{\mathcal{X}}\times{\mathcal{U}}\rightarrow{\mathbb{R}}$, and $m$ initial states $X^{k}_{0}\in{\mathcal{X}},k\in[m]$, in some fashion (deterministic or random). Note that the $m$ initial states need not be distinct.
  2. 2.
    At time $t$, for each action index $k\in[m]$, with current state $X^{k}_{t}=x^{k}_{i}$, choose the current action as $U_{t}=u_{k}\in{\mathcal{U}}$, and let the Markov process run for one time step. Observe the resulting next state $X^{k}_{t+1}=x^{k}_{j}$. Then update function $Q_{t}$ as follows, once for each $k\in[m]$:
    $Q_{t+1}(x^{k}_{i},u_{k})=\left\{\begin{array}[]{ll}Q_{t}(x^{k}_{i},u_{k})+\alpha_{t,i,k}[R(x_{i},u_{k})+\gamma V_{t}(x^{k}_{j})-Q_{t}(x^{k}_{i},u_{k})],&\mbox{if }x_{s}=x^{k}_{i},\\ Q_{t}(x^{k}_{s},u_{k}),&\mbox{if }x^{k}_{s}\neq x^{k}_{i}.\end{array}\right.$ (5.12)
    where
    $V_{t}(x^{k}_{j})=\max_{w_{l}\in{\mathcal{U}}}Q_{t}(x^{k}_{j},w_{l}).$ (5.13)
    Here $\alpha_{t,i,k}$ equals $\beta_{t}$ for all $i,k$ if a global clock is used, and equals
    $\alpha_{t,i,k}=\sum_{\tau=0}^{t}I_{\{X^{k}_{t}=x_{i}\}}$ (5.14)
    if a local clock is used.
  3. 3.
    Repeat.
Remark: Note that $m$ different simulations are being run in parallel, and that in the $k$-th simulation, the next action $U_{t}$ is always chosen as $u_{k}$. Hence, at each instant of time $t$, exactly $m$ components of $Q(\cdot,\cdot)$ (viewed as an $n\times m$ matrix) are updated, namely the $(X^{k}_{t},u_{k})$ component, for each $k\in[m]$.
Now, by fitting this algorithm into the framework of Theorems 4.2 and 4.3, we can prove the following general result.
Theorem 5.2.
Suppose that each matrix $A^{u_{k}}$ is irreducible, and that the step size sequence $\{\beta_{t}\}$ satisfies the Robbins-Monro conditions
$\sum_{t=0}^{\infty}\beta_{t}^{2}<\infty,\sum_{t=0}^{\infty}\beta_{t}=\infty.$ (5.15)
With this assumption, we have the following:
  1. 1.
    If a local clock is used, then $Q_{t}$ converges almost surely to $Q^{*}$.
  2. 2.
    If a global clock is used, and $\{\beta_{t}\}$ is nonincreasing, then $Q_{t}$ converges almost surely to $Q^{*}$.
We give only a sketch of the proof. The quantity being multiplied by $\alpha_{t,i,k}$ in (5.12) is $R(x_{i},u_{k})+\gamma V_{t}(x_{j})-Q_{t}(x_{i},u_{k})$, where $x_{j}$ is a randomly generated next state corresponding to the action $u_{k}$. The key observation, which is apparently made for the first time in [38], is this:
$E\left[\max_{w_{l}\in{\mathcal{U}}}Q_{t}(X_{t+1},w_{l})|(X_{t},U_{t})=(x_{i},u_{k})\right]=\sum_{j=1}^{n}a^{u_{k}}_{ij}\max_{w_{l}\in{\mathcal{U}}}Q_{t}(x_{j},w_{l}).$
Hence the quantity
$V_{t}(x_{j})=\max_{w_{l}\in{\mathcal{U}}}Q_{t}(x_{j},w_{l})$
is an unbiased estimate of
$\max_{w_{l}\in{\mathcal{U}}}Q_{t}(X_{t+1},w_{l})|(X_{t},U_{t}).$
Hence, if we define the measurement error as
$\xi_{t,i,k}=R(x_{i},u_{k})+\gamma\sum_{j=1}^{n}a_{ij}V_{t}(x_{j})-Q_{t}(x_{i},u_{k})-\{R(x_{i},u_{k})+\gamma V_{t}(x_{j})-Q_{t}(x_{i},u_{k})\},$
then $\xi_{t,i,k}$ has zero conditional mean and satisfies (4.8) with $\mu_{t}=0$ for all $t$. Also, the conditional variance of $\xi_{t,i,k}$ grows quadratically as a function of the Euclidean norm-squared of $Q(\cdot,\cdot)$, so that (4.9) is also satisfied for some constant $M$ that does not depend on $t$. Hence the batch $Q$-learning algorithm proposed here is just an implementation of (1.2) for finding a fixed point of the contractive map $F$ defined in (5.7). The difference between the present Theorem 5.2 and previous results is this: In earlier work, it is just assumed that the Robbins-Monro type of conditions (5.10) and (5.11) hold. Out of these, (5.11) is easy, if the step size sequence is square-summable. However, the divergence condition (5.10) is not so easy to establish. In the present case, because there are $m$ simulations running in parallel, (5.15) (together with a monotonicity assumption if a global clock is used) suffice to establish the convergence. These conclusions are a ready consequence of the results in Section 4.
In earlier papers, the satisfaction of (5.10) is sought to be achieved by assuming that every policy $\pi$ leads to an irreducible Markov process. It is shown in [39] that the following problem is NP-hard: Given an MDP, determine whether every policy $\pi$ results in a Markov process that is a unichain, that is, consists of a single set of recurrent states with the associated state transition matrix being irreducible, plus possibly some transient states. Our problem is slightly different, because we don’t permit any transient states. Nevertheless, this problem is also likely to be very difficult. In contrast, it is assumed in Theorem 5.2 that every action leads to an an irreducible Markov process. In other words, the assumption is that the $m$ different matrices $A^{u_{k}},k\in[m]$ correspond to irreducible Markov processes. This is assumption is very easily verified through algebraic methods. Thus Theorem 5.2 represents a substantial improvement over known results, as the conditions in (5.12) are automatically satisfied.
6 Conclusions and Problems for Future Research
In this paper, we have reviewed some results on the convergence of the Stochastic Gradient method from [17]. Then we analyzed the convergence of “intermittently updated” processes of the form (4.1). For this formulation, we derived sufficient conditions for convergence, as well as bounds on the rate of convergence. Building on this, we derived both sufficient conditions for convergence, and bounds on the rate of convergence, for the full BASA formulation of (1.2). Next, we applied these results to derive sufficient conditions for the convergence of the well-known $Q$-learning algorithm of Reinforcement Learning. In existing analysis, the convergence is established assuming that an analog of the standard Robbins-Monro conditions holds. However, verifying these conditions is not trivial; indeed, this problem is very similar to an NP-hard problem; see Section 5 for details. In this paper, by introducing a batch version of $Q$-learning, we were able to do away with these unverifiable conditions, and replace them by simple algebraic conditions.
There are several interesting problems thrown up by the analysis here. To our knowledge, our paper is the first to provide explicit estimates of the rates of convergence for BASA. Despite this, we have not provided rates of convergence for the $Q$-learning algorithm using our methodology. (The paper [41] uses an entirely different approach, and the bounds derived therein are not directly comparable to our bounds.) A related issue is that of “Markovian” stochastic approximation, in which the update process is the sample path of an irreducible Markov process. It would be worthwhile to examine whether the present approach can handle Markovian SA as well.